### v1

# thanks: Present Address: Imec, Kapeldreef 75, 3001 Leuven, Belgium

Abstract
We study the frustration pattern of a square lattice with in-situ fabricated Nb-Pt-Nb four-terminal Josephson junctions. The four-terminal geometry gives rise to a checker board pattern of alternating fluxes $f$, $f^{\prime}$ piercing the plaquettes, which stabilizes the Berezinskii–Kosterlitz–Thouless transition even at irrational flux quanta per unit cell, due to an unequal repartition of integer flux sum $f+f^{\prime}$ into alternating plaquettes. This type of frustrated frustration manifests as a beating pattern of the dc resistance, with state configurations at the resistance dips gradually changing between the conventional zero-flux and half-flux states. Hence, the four-terminal Josephson junctions array offers a promising platform to study previously unexplored flux and vortex configurations, and provides an estimate on the spatial expansion of the four-terminal Josephson junctions central weak link area.
Since the 1980’s, arrays of Josephson junctions are being investigated [1] and a broad base of knowledge about the physics of these arrays has been accumulated over the years [2, 3, 4, 5, 6, 7]. Recent findings in the field include the engineering of energy-phase relations with arrays [8], a deeper understanding of the vortex-lattice states in arrays [9], the demonstration of giant fractional Shapiro steps in anisotropic arrays [10], and the creation of arrays made of superconducting islands on a normal-conducting weak link material [10, 11, 12, 13].
Typically, the Josephson junction arrays are formed by two-terminal junctions. Recently, however, multi-terminal Josephson junctions received increasing attention [14, 15, 16]. In general, a multi-terminal Josephson junction is defined by multiple superconducting leads being connected by a central weak link region [14, 15]. Various weak-link materials can be used for these multi-terminal Josephson junctions. Using superconductor-semiconductor based multi-terminal Josephson junctions, phase control over Andreev bound states [17], control of the diode effect by a gate [18] or by a magnetic flux [19], as well as access to the single-mode conductance regime [16, 20] could be demonstrated. Multi-terminal Josephson junctions with topological insulator as weak link material have been realized [21] and a superconducting diode effect has been observed here as well [22, 23, 24]. Topological insulator based multi-terminal Josephson junctions are considered as key element in various Majorana fermion braiding architectures [25, 26, 27]. More generally, a multi-terminal Josephson junction is predicted to host topological states without requiring any topological material [15].
As illustrated in Fig. 1 a), by means of a multi-terminal junction with four terminals a novel type of Josephson junction array can be created.
Figure 1: a) Schematic illustration of four-terminal Josephson junctions forming the array by connecting their superconducting arms. Each blue circle represents a 4TJJ with its four superconducting Nb fingers being connected by a central Pt normal conducting area (see inset). The superconducting fingers are connected to neighboring 4TJJs, forming a square lattice together. b) Schematics of the Josephson junction array where the 4TJJs are described by four interconnected two-terminal Josephson junctions (see dashed square) following the approach of Graziano et al. [28]. Upon application of a magnetic field each large unit cell area ($A$) of the array is penetrated by a magnetic flux $\Phi$. In addition, the weak link area ($A^{\prime}$) of the 4TJJ (blue colored area) is penetrated by a flux $\Phi^{\prime}$. The sum of both areas equals $A_{\text{uc}}$. c) Theoretical representation of the array shown in b) as a 2TJJ array with a square unit cell and alternating frustrations $f=\Phi/\Phi_{0}$ (grey) and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$ (blue). d) False-color scanning electron microscope image of a 5$\times$5 four-terminal Josephson junction array. The stencil mask (green) was only partially removed, enabling a view on the mask and the formed array below. The platinum (blue) is deposited under rotation and, thus, covers a larger area. The superconducting Nb contacts (red) form the superconducting electrodes of a 4TJJ where their ends meet in the central Pt weak link area. The unit cell depicts area $A_{\text{uc}}$. e) Scanning electron micrograph of the 30$\times$30 4TJJ array presented in this study. Due to the shadowing effects deforming the leads left and right from the array, there are a few four-terminal junctions less in the array.
In contrast to conventional Josephson junction arrays, where a two-terminal Josephson junction (2TJJ) is placed in each arm forming a plaquette of the array, here a four-terminal Josephson junction (4TJJ) is situated in each corner of the plaquette. We aim to use the overlap between the two fields, arrays and multi-terminal Josephson junctions, to generate new knowledge for both of them.
We investigate Josephson junction arrays with a 4TJJ in each corner of a plaquette forming the array. Our 4TJJs are based on superconductor/normal conductor structure, using a metal as a weak link between the superconducting electrodes [1, 2, 29, 10, 30, 13]. More specifically, we use Nb for the four closely spaced superconducting electrodes connected by a metallic Pt weak link. The manufacturing process of the 4TJJ array is based on a stencil lithography process using molecular beam epitaxy with a high device yield [31]. The shadow evaporation based fabrication scheme employed here also ensures ultra-clean interfaces between the Nb and Pt layers. The periodic features in the magnetoresistance of a $30\times 30$ 4TJJ array are investigated in detail taking into account the magnetic flux penetrating each plaquette and the flux in the weak link area of the 4TJJ. The experimental results are supported by numerical simulations based on the resistively capacitively shunted junction (RCSJ) model, where the 4TJJ at each crossing point of the array is modelled by four interconnected two-terminal junctions, as illustrated in Fig 1 b).
The stencil lithography fabrication process of the 4TJJ array with molecular beam epitaxy is described in the following. As a first step, a Si(111) substrate is covered by a 300-nm-thick SiO2 layer followed by a 100-nm-thick Si3N4 layer. Subsequently, the stencil mask is patterned into the Si3N4 layer by electron beam lithography and reactive ion etching. As a next step, the stencil mask is under-etched by hydrofluoric acid, removing the SiO2 layer below and revealing the Si(111) substrate surface. The stencil mask can be seen in Fig. 1 d).
After inserting the pre-patterned substrate into an ultra high vacuum growth chamber, a 10-nm-thick Al2O3 interlayer is first deposited under rotation. The rotation ensures that the junction array area is completely covered by Al2O3. Next, a 20-nm-thick Pt layer is deposited under rotation. Here, the deposition at different angles due to the substrate rotation results in the deposition of Pt connecting the openings of the stencil mask. Finally, a 50-nm-thick superconducting Nb layer is deposited without rotation, resulting in a Nb electrode pattern defined by the opening of the stencil mask. As a last step, a 5-nm-thick Al2O3 capping layer is deposited under rotation. In Fig. 1 d) a scanning electron microscope image of the final device structure is shown. The four-terminal Josephson junction consists of four superconducting Nb islands connected at their tips by a central Pt weak link area. After the material deposition, the stencil mask is removed by applying ultrasonic power. The square unit cell of the array formed by a 4TJJ in each corner is indicated in Fig. 1 d). To prevent screening of magnetic fluxes by screening currents in superconducting material surrounding the array, the Nb layer outside the device is removed by reactive ion etching. Figure 1 e) shows a scanning electron microscope image of the 30$\times$30 4TJJ array investigated in this study. In addition to the device presented in the main text, we measured an identical array device and a reference two-terminal Josephson junction, all fabricated on the same substrate during the same fabrication run, presented in the supplementary material.
The measurements have been performed in a dilution refrigerator with a base temperature of 70 mK equipped with a superconducting magnet. For measurements with magnetic fields below 80 mT, an external high-resolution current source was used to supply the bias current of the magnet. The voltage signal, from which the resistance ($R_{\text{dc}}$) and differential resistance ($dV/dI$) has been calculated, was measured with a quasi 4-point setup by applying a dc bias to the array. In additional measurements, the dc signal was superimposed with a smaller AC signal in order to measure the differential resistance using the lock-in technique (see supplementary material).
Magnetotransport measurements have been performed on a $30\times 30$ 4TJJ array at a temperature of 80 mK. At zero magnetic field, the array shows a critical current of $I_{\mathrm{c}}=57\;\mu$A and, close to the superconducting regime, the device has a differential resistance of around 5.5 $\Omega$ (see supplementary material).
The differential resistance-bias current-magnetic field diagram depicted in Fig. 2 a) shows periodic oscillations of the critical current with magnetic field. When applying a direct current of 30 $\mu$A through the device, its resistance oscillates with the same periodicity in magnetic field (cf. Fig. 2 b), which was determined to be 6.25 mT by a fast Fourier transform (see supplementary material). In addition, a device- and array-independent (present in the reference 2TJJ and both arrays) magnetic hysteresis of the resistance pattern was measured, which is not further considered in the main text (see supplementary material).
[…]
The resistance signal is similar to what, e.g. Rzchowski et al. [2] measured, however, the expected frustration pattern for rational values of $f$ is missing. As can be seen in Fig. 2 c), at certain magnetic field around ±$\pm 5\;\Phi_{0}/A_{\mathrm{uc}}$, the oscillations first disappear and then reappear when further increasing the magnetic field. As will be explained below, both features, the different frustration pattern and the beating pattern, stem from the altered array lattice, which can be described by a conventional 2TJJ square lattice with checker board flux pattern. This gives rise to a superimposed beating pattern in addition to the flux quantum oscillations of the array unit cell. The periodicity of that beating pattern is approximately equal to the area ratio of the large array unit cell ($A$) and the 4TJJ central weak link region ($A^{\prime}$). The base offset of the resistance dips can be explained by the Fraunhofer effect of the 2TJJs between two individual contacts of a 4TJJ. The resistance oscillations can be seen to magnetic fields above ±$\pm 100\;$mT with, in total, 30 flux quantum oscillations ($30\;\Phi_{0}/A_{\mathrm{uc}}$).
Figure 2: Measurement data of the 30$\times$30 4TJJ array at 80 mK. a) Differential resistance as a function of bias current and magnetic field. A periodic oscillation of the critical current is clearly visible. b) Resistance as a function of magnetic field with an applied dc bias of 30 $\mu$A. The resistance oscillations correspond to a periodicity of 6.25 mT. At around ±$\pm 5\;\Phi_{0}/A_{\mathrm{uc}}$, the resistance oscillations are damped. c) Resistance oscillations under magnetic field up to ±$\pm 140$ mT with an applied dc bias of 30 $\mu$A. For magnetic fields higher than the point of damping, the oscillations stay present to above $\pm$100 mT.
To explain the measured behavior of the 4TJJ array, a corresponding theoretical model is introduced in the following. The experimental device is schematically depicted in Fig. 1 a). Its square unit cell has one 4TJJ in each corner. The junctions are connected via their superconducting Nb contacts and every junction has a central Pt weak link region. Based on the RCSJ network model for multi-terminal Josephson junctions [28], the array can be described as shown in Fig. 1 b) with four 2TJJs creating one 4TJJ (see also supplementary material). This introduces a second lattice of areas pierced by magnetic flux, the central weak link region of the 4TJJs.
The multi-terminal array model as shown schematically in Figs. 1 a) and b) can be conveniently represented in terms of an ordinary square lattice junction model [see Fig. 1 c)], with the following important difference to previous theoretical and experimental studies: instead of all the plaquettes being pierced by the same flux, we get an alternating (checker board) flux pattern ($f=\Phi/\Phi_{0}$ and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$). We capture the dynamics of this system by means of a classical RCSJ model approach, where the equations of motion for the superconducting phases $\phi_{j}$ at nodes $j=(j_{x},j_{y})$ follow as usual from the Kirchhoff laws (for more details see supplementary material).
Including a source and drain contact on two sides of the lattice, the RCSJ model allows for a direct computation of the dc resistance $R_{\text{dc}}$ as a function of the magnetic field. We do so by an explicit numerical evaluation of the classical RCSJ equations of motion. In addition we track the state of the system by a quantity related to the more commonly considered vorticity [4]: for a given plaquette, we calculate the counter clockwise supercurrent in the loop defining the plaquette (where for two neighbouring circuit nodes $\phi_{i},\phi_{j}$, the supercurrent is given as $I_{c}\sin(\phi_{i}-\phi_{j}+a_{ij})$ with the equivalent of the Peierl’s phase $a_{ij}$ capturing the magnetic flux), see also Fig. 3 a). This loop current captures the vortex structure qualitatively quite well, see, e.g., the loop current texture of the regular square lattice system $f^{\prime}=f$ for $f=1/2,1/3$ in Figs. 3 b) and c). Contrary to the vorticity however, the loop current does not return an integer number. This property will turn out to be useful in a moment, to describe stable states for the checker board flux system, $f^{\prime}\neq f$.
Figure 3: Results of the theoretical analysis. a) Checker board arrangement of alternating fluxes $f,f^{\prime}$, and depiction of superconducting loop currents (inset within green circle) as an alternative quantity to represent the vortex configuration. While loop currents do not provide an integer-valued quantity (contrary to vorticity), the ordinary $f=1/2$ (b) and $f=1/3$ (c) states are nontheless clearly visible by structures repeating either over $2$ or $3$ plaquettes, respectively. The dots indicate clockwise (blue) or counterclockwise (red) going currents and their size represents the magnitude of the loop current (relative linear scale). For the checker board ($f\neq f^{\prime}$), special points are those where $f+f^{\prime}$ adds up to an integer (d). At these points, the dc resistance (e) experiences superconducting dips and exhibits a beating pattern with large periodicity, consistent with experimental data. The current loop texture for the stationary states (no bias current) at the respective dips is shown in panels (f-i), which differ markedly from the usual stable (or semistable) states of conventional lattices with $f=f^{\prime}$ (b,c).
That being said, we explicitly compute $R_{\text{dc}}$ as a function of $f$ for a square lattice with $6\times 6$ nodes  [^1] The theory modelling occurs with a smaller system compared to the actual experimental setup. We do so simply to reduce the computational effort; the qualitative features of the theory model do not depend strongly on system size. for $f/f^{\prime}=5.4$, yielding Fig. 3 e). The central finding is this: the dc resistance experiences superconducting dips whenever $f=-f^{\prime}$ modulus $1$, see Fig. 3 d). Consequently, while the flux through a single plaquette is far from being zero at a given dip (or an integer of the flux quantum), the sum of two neighbouring plaquettes $f+f^{\prime}$ is. Denoting the ratio $f^{\prime}/f=\beta$, these dips occur for $f_{m}=m/(1+\beta)$ with $m\in\mathbb{Z}$. Since $\beta$ is given by the area ratio (see discussion of the experimental section above), $\beta$ is in general an irrational number, such that the dips in $R_{\text{dc}}$ likewise occur at irrational values, $f_{m}\notin\mathbb{Q}$. This is in stark contrast to regular square lattices commonly considered in previous works ($\beta=1$), where $R_{\text{dc}}$ dips at integer $f$, with weaker side dips for rational $f$ with low denominator. In addition to the dips repeating with period $p_{\text{dips}}=1/(\beta+1)$ in $f$ space, there is an overall beating pattern due to the fact that there repeatedly occur configurations where both $f_{m}$ and $\beta f_{m}$ (value of $f^{\prime}$ at the dip) return to a value close to zero (up to modulus $1$), where we approximately recover the state at zero applied field. The period for that pattern is $p_{\text{beating}}\approx 1/\beta$, where, due to $\beta\ll 1$ (large discrepancy between two relevant areas in experimental setup) $p_{\text{beating}}\gg p_{\text{dip}}$. This finding agrees very well with the experimental data shown in Figs. 2 b) and c).
When converting the frustration dip periodicity into magnetic field dip periodicity $p_{\text{dips,B}}$ (6.25 mT) with $1/(\beta+1)=p_{\text{dips,B}}\cdot A/\Phi_{0}$ and simplifying, we find $\Phi_{0}/p_{\text{dips,B}}=A^{\prime}+A=A_{\text{uc}}$. This result agrees with Fig. 2 b). In case that $A^{\prime}$ and $A$ are unknown, one can use the the difference in periodicities $p_{\text{beating}}-p_{\text{dips}}\approx 1/\beta-1/(\beta+1)$ to estimate $\beta$. Here, a value of $\beta\approx 0.092$ was determined (see supplementary material). The area of the central weak link region is given by $A^{\prime}=A\beta=(A_{\text{uc}}-A^{\prime})\beta$, resulting in
$A^{\prime}=A_{\text{uc}}\frac{\beta}{1+\beta}.$ (1)
With this, the central weak link region of the 4TJJs is $A^{\prime}\approx(165\;\text{nm})^{2}$. This is compatible with the 4TJJ geometry (see supplementary material).
The only feature that is not captured is the base offset of the resistance dips that occurs for some finite magnetic fields (the fact that the $R_{\text{dc}}$ dips return to a finite value instead of zero). This extra phenomenology can however be explained very easily: the above deployed RCSJ model assumes critical currents for the individual junctions that are independent of the magnetic field. In reality, given the estimated areas of the junctions, there is a finite onset of a Fraunhofer pattern, reducing the critical current at stronger magnetic fields (Fig. 2 a). When biasing the array with a constant current and ramping the magnetic field, this results in the finite-value- instead of zero-resistance dips.
Ultimately, the checker board arrangement of alternating fluxes ($f,f^{\prime}$) provides a unique opportunity to stabilize the superconducting (insulating) phase of the Berezinskii–Kosterlitz–Thouless (BKT) transition for incommensurate (irrational) flux quanta, since two neighbouring plaquettes can be tuned to values that cancel (up to modulus $1$) over a unit cell containing two plaquettes. The insulating phase of the vortex BKT transition corresponds to a superconducting phase in terms of the transport properties (which is why we refer to this phase as the superconducting phase for the remainder of this work). The special feature of stabilizing the superconducting phase for incommensurate (irrational) flux quanta motivates us to study the stationary system state (at zero bias current) at the values $f_{m}$ where $R_{\text{dc}}=0$ (suppressed vortex motion). Due to the irrational values for $f,f^{\prime}$, the (integer-valued) vorticity is not well-suited for such an analysis. This is why we instead compute the steady state configuration of the previously introduced loop currents in Figs. 3 f)-i). Even though $f,f^{\prime}\notin\mathbb{Q}$, the unit cell of the stationary state is always of the size of two plaquettes, due to $f+f^{\prime}$ cancelling – in contrast to, e.g., the regular $f=f^{\prime}=1/3$ state, which repeats over three plaquettes. In fact, the loop current texture of the respective stationary states gradually changes between the regular $f=f^{\prime}=0$ state (when both $f$ and $f^{\prime}$ are roughly integer) and the regular $f=f^{\prime}=1/2$ state (note that for half-integer flux quanta, $1/2$ is equivalent to $-1/2$, hence the checker board state with $f=-f^{\prime}$ is here indistinguishable from $f=+f^{\prime}$). The here presented work thus provides a glimpse into classes of states that have not been investigated so far. While the dips at the beginning and the end of a beating pattern (period $p_{\text{beating}}$) correspond very closely to the regular states $f=f^{\prime}$ integer, the dip in the middle of the beating period corresponds to $f=f^{\prime}=1/2$ – thus evidencing a regular frustration pattern. The dips in between provide states that defy the usual frustration picture as they occur at flux values far away from rational flux quanta. Therefore, we can think of the checker board arrangement with incommensurate ratio $\beta$ as a way to circumvent usual frustration.
Finally, the above finding can be cast into the more commonly used language of vortices and the BKT transition. The energy of the classical 2D Josephson junction array in terms of the number of vortices $v$ per plaquette is given by (see, e.g., Ref. [9] for details),
$H=\sum_{i,j}(v_{i}-f_{i})U_{ij}(v_{j}-f_{j})\ ,$ (2)
where the indices $i,j$ address the plaquette positions in 2 dimensions (i.e., both $i$ and $j$ implicitly encode a pair of $x$ and $y$ coordinates), and $U_{ij}$ represents the typical logarithmic long-range interaction. For regular square lattices $f_{i}=f$ (i.e., all plaquettes experience the same flux), the well-known BKT picture arises. At zero flux $f=0$, the zero vortex state $v_{j}=0$ (which is insulating in that no vortices are available to move) corresponds to the ground state, whereas states with finite vortex numbers or vortex pairs (conducting states) are at low temperatures inaccessible due to a finite energy gap, leading to the BKT transition. At finite flux, e.g., $f>0$, the system reacts in such a way that the new ground state can be to a good approximation described by a state with a finite vortex density of $\sim f$ vortices per plaquette, in order to screen the applied external flux, where (as is well-known) the propensity of the ground state to have ordered (and thus likewise insulating) vortex configurations or many freely moving vortex configurations depends on whether $f$ is a rational number or not.
For the here present checker board flux texture ($f_{i}=f,f^{\prime}$ for alternating plaquette positions) the situation is different in the following crucial aspect. The aforementioned special points of cancelling neighbouring fluxes $f^{\prime}=-f$ still have the zero vortex state $v_{j}=0$ as the gapped ground state because averaged over the whole system, no flux is applied, and hence no finite vortex density occurs due to screening. We therefore arrive again at the superconducting BKT phase, notably, at generally irrational values of $f$. We emphasise that these zero flux states are by no means trivial. Consider the special (rational) case of $f=1/2$ and $f^{\prime}=-1/2$, with the zero vortex ground state. With a simple unitary transformation (increasing the vortex number on every $-1/2$ plaquette by one), we map to the regular square lattice $f=f^{\prime}=1/2$ with the well-known ground state given by an alternating $0,1$ vortex configuration. Consequently, $f=-f^{\prime}=1/2$ with the zero vortex ground state and $f=f^{\prime}=1/2$ with alternating $0,1$ vortex ground state are equivalent. Now, the difference to the case where $f=-f^{\prime}$ but $f$ irrational (or in fact any rational or irrational number other than $f=0$ or $f=1/2$) is that the correspondence to the regular lattice $f=f^{\prime}$ is no longer exact (as there exists no unitary mapping due to $v_{i}\in\mathbb{Z}$). For values close to $1/2$, there still exists a unitary map to a system approximately equivalent to $f=f^{\prime}=1/2$, this correspondence grows weaker and weaker as we move away from $1/2$, which at the value of $1/4$ or lower leads to a cross over, where the checker board system with $f=-f^{\prime}$ more closely imitates the regular $f=f^{\prime}=0$ system instead of $f=f^{\prime}=1/2$. This constitutes a vortex perspective which is ultimately equivalent to the above analysis in terms of superconducting loop current configurations changing from the $f=f^{\prime}=0$ to the $f=f^{\prime}=1/2$ ground state [see also Figs. 3 f)-i) and their discussion above].
We demonstrated an array made with 4TJJs. For this we introduced an in-situ fabrication technique for array fabrication. A frustration pattern was measured which differs from the expectation for ordinary 2TJJ arrays. The difference could be explained by theoretically introducing a checker board lattice with two alternating flux patterns $f$ and $f^{\prime}$. The periodicity of the beating pattern can be connected to the area ratios $A$/$A^{\prime}$. In particular, while dips in the dc resistance can be linked to the sum $f+f^{\prime}$ being integer, the in general irrational area ratio leads to an irrational repartition of this total flux into the alternating plaquettes. This allows for the stabilization of the superconducting BKT phase at irrational fractions of flux penetrating individual plaquettes – a feature we termed frustrated frustration. This feature allows us to estimate the spatial expansion of the 4TJJ central weak link region, the result of which is compatible with the junction geometry. Overall, the here considered setup opens up a previously unexplored class of array system with alternating flux textures. Given that detailed understanding of vortex lattice structures and their dynamics are still an open question for regular lattices with a single flux $f$ [9], the here introduced alternating $f,f^{\prime}$ lattice will likely lead to a rich catalogue of open research directions of its own, the full extent of which we could not possibly anticipate at this stage. As an example of a practical near-term question, we expect that at lower critical currents it might be perceivable to measure signatures related to rational (instead of integer) values of the sum $f+f^{\prime}$, leading to substructures within the beating pattern profile. Also, non-local (e.g., multi-terminal) transport on such arrays, as well as their Shapiro response, could reveal further information about any nontrivial vortex configuration or motion.
We thank Herbert Kertz for technical assistance. All samples have been prepared at the Helmholtz Nano Facility [33]. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769, by the German Federal Ministry of Education and Research (BMBF) via the Quantum Futur project ‘MajoranaChips’ (Grant No. 13N15264), as well as the Bavarian Ministry of Economic Affairs, Regional Development and Energy within Bavaria’s High-Tech Agenda Project ”Bausteine für das Quantencomputing auf Basis topologischer Materialien mit experimentellen und theoretischen Ansätzen” (grant no. 07 02/686 58/1/21 1/22 2/23).
SUPPLEMENTAL NOTE 1: Second 30$\times$30 4TJJ Array
In addition to the 30$\times$30 4TJJ array presented in the main text, there was a second device on the same sample. It equals the first in every aspect. Here, we present the data of the second 4TJJ array (see Fig. S1).
The critical current oscillates with magnetic field in the same way as for the main text array. With a DC bias of 30 $\mu$A, that behavior leads to an oscillation of the resistance with magnetic field (see Fig. S4 b). Similar to the main text device, the frustrated frustration pattern is present. The damping of the oscillations of resistance and critical current can be reproduced.
At zero magnetic field and a temperature of 80 mK, the critical current is comparable to the main text device with $I_{\mathrm{c}}=59.75\;\mu$A (see Fig. S1 c and d). The differential resistance is lower with $R=1.75\;\Omega$ close to the maximum applied current of 100 $\mu$A.
Supplemental Figure S1: Data of the second 30$\times$30 four-terminal Josephson junction array. a) Scanning electron microscopy image of the device. b) Differential resistance-bias current-magnetic field diagram. The oscillations of the critical current are clearly visible. c) I-V measurement at 80 mK. d) Differential resistance calculated from the data in figure c).
SUPPLEMENTAL NOTE 2: Reference two-terminal Josephson junction
On the chip of the presented arrays, a reference two-terminal Josephson junction has been fabricated in the same way as the arrays (see figure S2). The weak link length was determined to be 92 nm and its width is 920 nm. Although only a width of 809 nm is indicated in figure S2, Pt and Nb material under the edge of the bottom Si3N4 layer needs to be taken into account. The angle between the substrate surface and niobium beam was 57.7. The 300 nm-thick-SiO2 layer right at the edge of the trench is etched away, just like under the mask itself. This leaves the 100 nm-thick-Si3N4 layer under-etched at the edge and gives additional substrate space for the niobium shadow to be deposited. The extent of the Nb shadow under the Si3N4 edge can be derived from the top shadow of image S2 (174 nm). The bottom shadow has a smaller distance from its edge than the top shadow because it is delimited by the lower edge of the Si3N4 layer (instead of the upper edge). That means that the bottom shadow will be $100\;\text{nm}/\text{tan}(57.7^{\circ})=63\;\text{nm}$ shorter than the top shadow. This results in a total bottom shadow of 111 nm, which has already been added to the width specification of this Josephson junction above (920 nm).
In figure S2, the junction is shifted with respect to the stencil mask (by 42.2) because latter was rotated against the niobium beam during deposition.
Supplemental Figure S2: Scanning electron microscope images of the reference two-terminal Josephson junction.
Figure S3 shows the voltage and differential resistance dependent on the current. The device has a critical current of 56.51 $\mu$A and a differential resistance of 1.21 $\Omega$. The Josephson energy of the junction is $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=116$ meV. Using a relative permittivity of one (because platinum is a metal) and the geometrical values given above (length of a SC contact is 600 $\mu$m), the coplanar junction capacitance is $C=5.6\cdot 10^{-17}\;$.
[…]
SUPPLEMENTAL NOTE 3: Magnetic hysteresis of the resistance pattern
A magnetic hyteresis has been observed when measuring the resistance with magnetic field. The resistance oscillation pattern is shifted for different sweep directions of the magnetic field (see Fig. S4). The shift has been observed in all devices measured: In the arrays as well as in the two-terminal Josephson junction. Therefore, we rule out that the effect is device-specific. A possible reason is the screening of magnetic field by present Abrikosov vortices.
(a) Main text 4TJJ array.
(b) Second 4TJJ array.
(c) Reference 2TJJ.
Supplemental Figure S4: Resistance dependent on magnetic field for: (a) the main text 4TJJ array, (b) the second 4TJJ array, (c) and the reference two-terminal Josephson junction. The blue lines indicate a magnetic sweep direction from negative to positive and the orange lines from positive to negative. All devices show hysteretic behavior with magnetic field. The two upper resistance oscillations stem from the frustration pattern of the array, the lower resistance oscillation stems from the superconducting quantum interference of the two-terminal Josephson junction.
SUPPLEMENTAL NOTE 4: Determination of the resistance oscillation period
For the determination of the oscillation frequencies, fast Fourier transforms have been applied to the data in Fig. 3 c) and d) of the main text (see Fig. S5). The frequency spectra of both data curves match. The dominant frequency (period) is 0.16 1/mT (6.25 mT). The period of the beating pattern cannot be seen in the spectrum, probably because it is too close to the zero peak.
Supplemental Figure S5: Fast Fourier transforms of the data curves in main text Fig. 3 c) (fine sweep) and d) (wide range). The dominant frequency (period) is 0.16 1/mT (6.25 mT). The other peaks of the spectrum are higher harmonics of that frequency.
SUPPLEMENTAL NOTE 5: Description of the array device
This section gives a detailed description of the array device in the main text, including geometrical properties of the array lattice (see Fig. S6). According to figure S6 a), the unit cell of the array is a square of size $A_{\text{UC}}=573\;\text{nm}\cdot 567\;\text{nm}=(570\;\text{nm})^{2}$. The geometry of the weak link area between two neighboring leads of one 4TJJ (see Fig. S6 b) is described, on average, by the following dimensions: The width of the weak link is 129 nm and its length is 97 nm. The 4TJJ leads have a length of 471 nm.
During fabrication of the device, a trench is etched into the SiO2 (300 nm) / Si2N3 (100 nm) layer, down to the Si(111) substrate. On the substrate, the device is fabricated and, afterwards, the whole chip is covered with superconducting material. To minimize disturbance effects, the superconductor around the device is etched away as a last fabrication step, leaving an edge of superconductor of approximately 1 µm around the trench behind. To prevent this closed loop of superconductor from generating screening currents, the edge is disconnected close to the ends of the device contacts (see figure S6 c-d).
Supplemental Figure S6: Scanning electron microscope images of the main text array. a): Description of the lattice geometry. b): Geometry of one 4TJJ in the array. c): Etching of the superconducting edge going around the trench of the device (left). The etchings are located close to the contact endings. A bond wire is visible in the middle. d): Close-up of a corresponding etching. The right side is the contact trench on the Si(111) substrate. The left side is the SiO2 (300 nm) / Si2N3 (100 nm) layer around the device.
SUPPLEMENTAL NOTE 6: Characteristic RCSJ values between two leads of a 4TJJ in the arrray
In this section, relevant parameters for the theoretical description of the array are derived from geometrical considerations of the prior section, as well as from measurement data of the reference two-terminal Josephson junction. We assume that critical current density and resistivity, derived in section 2, are geometry independent. We then use these values, together with the 4TJJ geometry to derive capacitance, resistivity, and critical current between two leads of one 4TJJ.
The coplanar capacitance is calculated using the geometrical values given above. The relative permeability $\epsilon_{r}$ is assumed to be one, since the weak link is a metal. This results in a capacitance of $C=2.729\cdot 10^{-18}$ F. The critical current between two 4TJJ leads is $I_{\mathrm{C}}=7.9$ µA. This results in a Josephson energy of $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=16.2$ meV. The resistance is equal to $R=9.1\;\Omega$.
SUPPLEMENTAL NOTE 7: Additional measurements of the main text 30$\times$30 4TJJ array
The current-voltage characteristics of the device presented in the main text can be seen in Fig. S7. The device has a critical current of $57\;\mu$A and a differential resistance of around 5.5 $\Omega$ close to the maximum applied current of 100 $\mu$A.
Supplemental Figure S7: IV measurements at 80 mK. a) The DC IV curve with a critical current of $I_{\mathrm{c}}=57\;\mu$A. b) AC differential resistance measured with the lock-in technique. The AC bias was 200 nA.
Additional measurements have been performed on the main text 30$\times$30 array device. Figure S8 shows temperature dependent IV curves. The declining critical current with increasing temperature is clearly visible.
Supplemental Figure S8: IV measurements dependent from 70 mK to 900 mK. Each measurement is shifted by -0.01 $\mu$V from the one before. b): Zoom in of a).
Figure S9 shows the same resistance oscillations with magnetic field as before but, here, they were measured with the AC lock-in technique and no DC bias. The resulting resistance is lower than with the DC measurement because the time-dependent sinusoidal signal includes the superconducting state of the device as well. When comparing figure S9 a) with figure S4 a), one can see that the oscillations are identical in magnetic field, although the peak pattern is more symmetrical with the AC method.
(a) $I_{\mathrm{AC}}=30\;\mu$A.
(b) $I_{\mathrm{AC}}=40\;\mu$A.
Supplemental Figure S9: AC resistance measurement dependent on magnetic field for 30 $\mu$A (a) and 40 $\mu$A (b). The blue lines indicate a magnetic sweep direction from negative to positive and the orange lines from positive to negative.
Figure LABEL:fig:SUPP_MISC_RB_DC_I_T a) shows resistance measurements dependent on magnetic field and DC current. At higher currents, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
Figure LABEL:fig:SUPP_MISC_RB_DC_I_T b) shows resistance measurements dependent on magnetic field and temperature. At higher temperatures, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
Supplemental Figure S10: DC resistance measurements dependent on magnetic field. a): At currents between 20 $\mu$A and 60 $\mu$A. b): At a DC bias of $I_{\mathrm{DC}}=30\;\mu$A and temperatures between 70 mK and 800 mK.
SUPPLEMENTAL NOTE 8: Details of the theoretical modelling
We here summarize the model calculations for the theoretical discussion of the main text. We describe the multiterminal Josephson junction array in the overdamped regime (discarding capacitive contributions), we get at each node $j$ the equation of motion,
$\frac{1}{2eR}\sum_{\langle j^{\prime}\rangle_{j}}\left(\dot{\phi}_{j}-\dot{\phi}_{j^{\prime}}\right)=-I_{c}\sum_{\langle j^{\prime}\rangle_{j}}\sin(\phi_{j}-\phi_{j^{\prime}}+a_{jj^{\prime}})+\sum_{\langle j^{\prime}\rangle_{j}}\eta_{jj^{\prime}}(t)\ ,$ (S1)
with shunt resistance $R$, critical current $I_{c}$, and the current noise term $\eta_{jj^{\prime}}$ due to thermal fluctuations from the resistor connecting nodes $j$ and $j^{\prime}$, where
$\langle\eta_{ii^{\prime}}(t)\eta_{jj^{\prime}}(t^{\prime})\rangle=\frac{2k_{B}T}{R}\delta_{ij}\delta_{i^{\prime}j^{\prime}}\delta(t-t^{\prime})\ .$ (S2)
The notation $\sum_{\langle j^{\prime}\rangle_{j}}$ indicates that the sum $j^{\prime}$ goes over nearest neighbours of $j$.
SUPPLEMENTAL NOTE 9: Estimation of the central weak link area
According to the theoretical discussion in the main text, the parameter $\beta$ and the frustration dip periodicities are defined as:
$\beta=\frac{f^{\prime}}{f}=\frac{A^{\prime}}{A}$ (S3)
$p_{\text{dips}}=\frac{1}{\beta+1}$ (S4)
$p_{\text{beating}}\approx\frac{1}{\beta}$ (S5)
With this, the difference of the two periodicities can be estimated:
$\Delta p\approx\frac{1}{\beta}-\frac{1}{\beta+1}$ (S6)
For this estimate the $\Delta p$ was determined by counting the number of resistance dips between the two beating points in figure 2 c). Ten dips ($\Delta p\approx 10$) can be seen between the two beating points. This yields $\beta\approx 0.092$ and, thus, an estimated weak link area of $A^{\prime}=A_{\text{UC}}\cdot\beta/(1+\beta)\approx(165\;\text{nm})^{2}$.
For comparison, such a weak link area is sketched on one 4TJJ in Fig. S11, assuming the weak link region to be a perfect circle with the corresponding area of $(165\;\text{nm})^{2}$. As can be seen, a shape of this size is compatible with the 4TJJ geometry.
Supplemental Figure S11: A 4TJJ of the main text array device presented in the main text with the estimated central weak link area sketched in green. Shape and position of the weak link area were chosen freely, its area equals the one estimated by the beating pattern, $(165\;\text{nm})^{2}$.

### v2

@@
 We study the frustration pattern of a square lattice with in-situ fabricated Nb-Pt-Nb four-terminal Josephson junctions.
-The four-terminal geometry gives rise to a checker board pattern of alternating fluxes $f$, $f^{\prime}$ piercing the plaquettes, which stabilizes the Berezinskii–Kosterlitz–Thouless transition even at irrational flux quanta per unit cell, due to an unequal repartition of integer flux sum $f+f^{\prime}$ into alternating plaquettes.
+The four-terminal geometry gives rise to a checker board pattern of alternating fluxes $f$, $f^{\prime}$ piercing the plaquettes, which stabilizes the Berezinskii–Kosterlitz–Thouless transition even at irrational flux quanta per plaquette, due to an unequal repartition of integer flux sum $f+f^{\prime}$ into alternating plaquettes.
 This type of frustrated frustration manifests as a beating pattern of the dc resistance, with state configurations at the resistance dips gradually changing between the conventional zero-flux and half-flux states.
-Hence, the four-terminal Josephson junctions array offers a promising platform to study previously unexplored flux and vortex configurations, and provides an estimate on the spatial expansion of the four-terminal Josephson junctions central weak link area.
-Since the 1980’s, arrays of Josephson junctions are being investigated [1] and a broad base of knowledge about the physics of these arrays has been accumulated over the years [2, 3, 4, 5, 6, 7].
-Recent findings in the field include the engineering of energy-phase relations with arrays [8], a deeper understanding of the vortex-lattice states in arrays [9], the demonstration of giant fractional Shapiro steps in anisotropic arrays [10], and the creation of arrays made of superconducting islands on a normal-conducting weak link material [10, 11, 12, 13].
-Typically, the Josephson junction arrays are formed by two-terminal junctions.
-Recently, however, multi-terminal Josephson junctions received increasing attention [14, 15, 16].
+Hence, the four-terminal Josephson junction array offers a promising platform to study previously unexplored flux and vortex configurations, and provides an estimate on the spatial expansion of the four-terminal Josephson junction central weak link area.
+Introduction.
+Arrays of Josephson junctions have been studied since the 1980s [1] and a broad base of knowledge about the physics of these arrays has been accumulated over the years [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
+Recent findings in the field include the engineering of energy-phase relations with arrays [20], a deeper understanding of the vortex-lattice states in arrays [2], the demonstration of giant fractional Shapiro steps in anisotropic arrays [22], and the creation of arrays made of superconducting islands on a normal-conducting weak link material [22, 23, 24, 25, 26].
+Typically, Josephson junction arrays are formed by two-terminal junctions.
+Recently, however, multi-terminal Josephson junctions received increasing attention [27, 28, 29, 30, 31].
 In general, a multi-terminal Josephson junction is defined by multiple superconducting leads being connected by a central weak link region [31, 28].
 Various weak-link materials can be used for these multi-terminal Josephson junctions.
-Using superconductor-semiconductor based multi-terminal Josephson junctions, phase control over Andreev bound states [17], control of the diode effect by a gate [18] or by a magnetic flux [19], as well as access to the single-mode conductance regime [16, 20] could be demonstrated.
-Multi-terminal Josephson junctions with topological insulator as weak link material have been realized [21] and a superconducting diode effect has been observed here as well [22, 23, 24].
+There is a broad spectrum of studies of devices with weak links such as semiconductors [32, 33, 34, 35, 36], graphene [29, 37, 38, 39], and topological insulators [40, 41, 42, 43].
 Topological insulator based multi-terminal Josephson junctions are considered as key element in various Majorana fermion braiding architectures [44, 45, 46].
 More generally, a multi-terminal Josephson junction is predicted to host topological states without requiring any topological material [28].
-As illustrated in Fig. 1 a), by means of a multi-terminal junction with four terminals a novel type of Josephson junction array can be created.
-Figure 1: a) Schematic illustration of four-terminal Josephson junctions forming the array by connecting their superconducting arms.
-Each blue circle represents a 4TJJ with its four superconducting Nb fingers being connected by a central Pt normal conducting area (see inset).
-The superconducting fingers are connected to neighboring 4TJJs, forming a square lattice together.
-b) Schematics of the Josephson junction array where the 4TJJs are described by four interconnected two-terminal Josephson junctions (see dashed square) following the approach of Graziano et al.
-[28].
-Upon application of a magnetic field each large unit cell area ($A$) of the array is penetrated by a magnetic flux $\Phi$.
-In addition, the weak link area ($A^{\prime}$) of the 4TJJ (blue colored area) is penetrated by a flux $\Phi^{\prime}$.
-The sum of both areas equals $A_{\text{uc}}$.
-c) Theoretical representation of the array shown in b) as a 2TJJ array with a square unit cell and alternating frustrations $f=\Phi/\Phi_{0}$ (grey) and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$ (blue).
-d) False-color scanning electron microscope image of a 5$\times$5 four-terminal Josephson junction array.
-The stencil mask (green) was only partially removed, enabling a view on the mask and the formed array below.
+We here report on the fabrication and study of a square array comprised of multi-terminal Josephson junctions, as illustrated in Fig. 1 a).
+Figure 1: a) Schematics of the Josephson junction array where each 4TJJ is described by four interconnected two-terminal Josephson junctions (see dashed circle) [29, 30, 37, 39].
+Their weak link material is platinum (see inset).
+The array is formed by connecting the superconducting arms of the 4TJJs.
+Upon application of a magnetic field each large plaquette area ($A$) of the array is penetrated by a magnetic flux $\Phi$ and the 4TJJ weak link area ($A^{\prime}$) by a flux $\Phi^{\prime}$.
+b) Theoretical representation of the array shown in a) as a 2TJJ array with square plaquettes having alternating frustrations $f=\Phi/\Phi_{0}$ (grey) and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$ (blue).
+c) False-color scanning electron micrograph of a $5\times 5$ four-terminal Josephson junction array with partially removed stencil mask.
 The platinum (blue) is deposited under rotation and, thus, covers a larger area.
-The superconducting Nb contacts (red) form the superconducting electrodes of a 4TJJ where their ends meet in the central Pt weak link area.
-The unit cell depicts area $A_{\text{uc}}$.
-e) Scanning electron micrograph of the 30$\times$30 4TJJ array presented in this study.
-Due to the shadowing effects deforming the leads left and right from the array, there are a few four-terminal junctions less in the array.
-In contrast to conventional Josephson junction arrays, where a two-terminal Josephson junction (2TJJ) is placed in each arm forming a plaquette of the array, here a four-terminal Josephson junction (4TJJ) is situated in each corner of the plaquette.
-We aim to use the overlap between the two fields, arrays and multi-terminal Josephson junctions, to generate new knowledge for both of them.
-We investigate Josephson junction arrays with a 4TJJ in each corner of a plaquette forming the array.
-Our 4TJJs are based on superconductor/normal conductor structure, using a metal as a weak link between the superconducting electrodes [1, 2, 29, 10, 30, 13].
+The niobium (red) forms the superconducting contacts.
+The dashed square represents unit cell area $A_{\text{uc}}$.
+d) Scanning electron micrograph of the $30\times 30$ 4TJJ array presented in this study.
+Due to the shadow evaporation process, the 4TJJ array is slightly deformed at the left and right ends of the array.
+In contrast to conventional Josephson junction arrays, where a two-terminal Josephson junction (2TJJ) is placed in each arm forming a plaquette of the array, here, a four-terminal Josephson junction (4TJJ) is located in each corner of the unit cell.
+As it turns out, this system has not a single flux parameter (as usual arrays do) but an alternating pattern of two fluxes.
+This allows us to find a stable superconducting phase for the array, even for irrational fluxes piercing the individual plaquettes.
+Our work thus connects vortex dynamics and the Berezinskii–Kosterlitz–Thouless (BKT) transition to the emerging field of incommensurable, quasiperiodic physics in solid state systems [47, 48, 49, 50, 51, 52].
+While the usual notion of quasiperiodic materials refers to lattices in real space, the incommensurability here is detectable in the space spanned by the applied flux.
+Experiments.
+Our 4TJJ array is based on a superconductor/normal conductor structure, using a metal as a weak link between the superconducting electrodes [1, 6, 27, 22, 53, 25].
 More specifically, we use Nb for the four closely spaced superconducting electrodes connected by a metallic Pt weak link.
-The manufacturing process of the 4TJJ array is based on a stencil lithography process using molecular beam epitaxy with a high device yield [31].
-The shadow evaporation based fabrication scheme employed here also ensures ultra-clean interfaces between the Nb and Pt layers.
-The periodic features in the magnetoresistance of a $30\times 30$ 4TJJ array are investigated in detail taking into account the magnetic flux penetrating each plaquette and the flux in the weak link area of the 4TJJ.
-The experimental results are supported by numerical simulations based on the resistively capacitively shunted junction (RCSJ) model, where the 4TJJ at each crossing point of the array is modelled by four interconnected two-terminal junctions, as illustrated in Fig 1 b).
-The stencil lithography fabrication process of the 4TJJ array with molecular beam epitaxy is described in the following.
+The manufacturing process of the 4TJJ array is based on a stencil lithography process using molecular beam epitaxy with a high device yield by ensuring ultra-clean interfaces between the Nb and Pt layers [54] (see Fig.
+1 and [55]).
+Figure 1 d) shows a scanning electron microscope image of the $30\times 30$ 4TJJ array investigated in this study.
+In addition to the device presented in the main text, we measured an identical array device and a reference two-terminal Josephson junction, all fabricated on the same substrate during the same fabrication run [55].
+Magnetotransport measurements have been performed on a $30\times 30$ 4TJJ array at a temperature of 80 mK.
+Without external magnetic field, the array shows a critical current of $I_{\mathrm{c}}=57\;\mu$A and, close to the superconducting regime, the device has a differential resistance of around 5.5 $\Omega$ [55].
+The differential resistance vs.
+bias current and out-of-plane magnetic field depicted in Fig.
+2 a) shows periodic oscillations of the critical current with magnetic field.
+When applying a fixed dc current of 30 $\mu$A through the device, its resistance oscillates with the same periodicity in magnetic field [cf.
+Fig.
+2 b)], which was determined to be 6.25 mT by a fast Fourier transform [55].
+In addition, a device- and array-independent magnetic hysteresis of the resistance pattern was measured, also present in the reference 2TJJ and both arrays, which is not discussed further in the main text (see Sec.
+V in the supplementary material).
+[…]
+[…]
+Figure 2: Measurement data of the $30\times 30$ 4TJJ array at 80 mK.
+a) Differential resistance as a function of bias current and magnetic field.
+A periodic oscillation of the critical current is clearly visible.
+b) Resistance as a function of magnetic field with an applied dc bias of 30 $\mu$A.
+The resistance oscillations correspond to a periodicity of 6.25 mT.
+At around ±$\pm 5\;f_{\mathrm{uc}}$, the resistance oscillations are damped.
+c) Resistance oscillations under magnetic field ranging to ±$\pm 140$ mT with an applied dc bias of 30 $\mu$A.
+In total, 30 flux quantum oscillations are present.
+Theoretical discussion.
+To explain the measured behavior of the 4TJJ array, we introduce a corresponding theoretical model.
+Based on the resistively capacitively shunted junction (RCSJ) network model for multi-terminal Josephson junctions [29, 30, 37, 39], the array can be described, as shown in Fig.
+1 a), with four 2TJJs creating one 4TJJ [55].
+This introduces a second lattice of areas pierced by magnetic flux, the central weak link region of the 4TJJs.
+The multi-terminal array model can therefore be conveniently represented in terms of an ordinary $J_{x}\times J_{y}$ square lattice junction model [where $J_{x}$ and $J_{y}$ are the horizontal and vertical number of superconducting nodes, see Fig.
+1 b)], with the following important difference to previous theoretical and experimental studies: instead of all the plaquettes being pierced by the same flux, we get an alternating (checker board) flux pattern ($f=\Phi/\Phi_{0}$ and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$).
+We deploy a classical RCSJ model approach, where the equations of motion for the superconducting phases at all nodes follow as usual from the Kirchhoff laws [55].
+We focus on the overdamped regime  [^1] Using the extracted values for individual junction capacitances, resistances, and critical currents, we estimate the Steward-McCumber parameter to be $\approx 5.4\times 10^{-6}$, well within the overdamped regime., neglecting the capacitive contributions to the equations of motion, and assume zero temperature.
+As will become apparent below, it is instructive to include a finite Fraunhofer effect [14] for the individual junctions in the lattice.
+Including a source and drain contact on two sides of the lattice (across which the bias current $I$ is applied), the RCSJ model allows for a direct computation of the dc resistance $R_{\text{dc}}$ as a function of the magnetic field.
+We do so by an explicit numerical evaluation of the classical RCSJ equations of motion  [^2] We use the NDSolve routine on Mathematica for explicit evaluation of the equations of mtion..
+The results of the simulation are summarized in Fig.
+3.
+For computational simplicity, the calculations were performed on a small array of $J_{x}=J_{y}=6$.
+It is instructive to first consider the special case $f,f^{\prime}\in\mathbb{Z}$ (equivalent to $f=f^{\prime}=0$).
+Here, the dc resistance can be found analytically (as both quantum and thermal phase slips are absent) [55],
+$\frac{R_{\text{dc}}}{R}=(J_{x}+1)\sqrt{\frac{1}{J_{y}^{2}}-\frac{I_{c}^{2}}{I^{2}}}\ ,$ (1)
+where $R$ and $I_{c}$ are the individual junction resistance and critical current, respectively.
+The array thus transitions from superconducting ($R_{\text{dc}}=0$) to resistive ($R_{\text{dc}}>0$) when the bias current exceeds $J_{y}I_{c}$.
+For a regular square lattice [$f=f^{\prime}$, see also Fig.
+3 b)] it is well known that the array generally leaves the superconducting regime for noninteger values of $f$, even though $I<J_{y}I_{c}$.
+For sufficiently low temperatures and current biases, the dc resistance nonetheless experiences dips at special rational values – known as the frustration pattern.
+Consider now the general checker board array, $f\neq f^{\prime}$ [left panel in Fig.
+3 b)], where we denote the (constant) ratio $\beta=f/f^{\prime}>1$.
+Here, there emerges a beating pattern, with the two characteristic periods $\beta>1$ and $\beta/(1+\beta)<1$ [see Figs.
+3 a) and c)].
+The larger period ($\beta$) is in general only approximate; e.g., for incommensurate $\beta$ the system can only approximately return to mutually integer $f,f^{\prime}$ – exhibiting a quasiperiodicity in flux space.
+The irrational nature of $\beta$ is however also highly relevant for the smaller period (which is exact): it separates the points where the sum of two neighbouring plaquette fluxes is zero, i.e., $f+f^{\prime}\in\mathbb{Z}$ [see top inset in Fig.
+3 a)].
+For the experimentally extracted value of $\beta\approx 10.9$ (cf.
+Fig.
+2[55]) the difference between the two frequencies is large, leading to a clearly visible beating pattern [Fig.
+3 c)].
+Once $\beta$ is known, the central weak link area of a 4TJJ can be determined via $A^{\prime}=A_{\text{uc}}/(1+\beta)$, leading to $A^{\prime}\approx(165\;\text{nm})^{2}$ [55].
+Crucially, for integer $f+f^{\prime}$, the repartition of the total flux into the two neighbouring plaquettes ($f$ and $f^{\prime}$) does not need to occur for special (integer or rational) value, since the ratio of the two fluxes, $\beta$, is in general incommesurate.
+Consequently, the system exhibits a stabilization of superconducting (BKT) phase (dips in $R_{\text{dc}}$) even for irrational $f,f^{\prime}$, a feature we choose to name ’frustrated frustration’.
+Figure 3: Results of the theoretical analysis.
+In all curves (a,c-g), the $x$-axis is $f$, and the $y$-axis is $R_{\text{dc}}/R$.
+(a) $R_{\text{dc}}$ as a function of $f$ for the checker board model with $f/f^{\prime}=10.9$ and $I/I_{c}=5.0$.
+(b) Checker board ($f\neq f^{\prime}$) versus regular ($f=f^{\prime}$) lattice models.
+(c) Zoomed out version of (a) showing the beating pattern.
+(d) $R_{\text{dc}}$ for the regular lattice model $f=f^{\prime}$, with all other parameters the same as in (a,c).
+(e,f) $R_{\text{dc}}$ including Fraunhofer effect (red curves, see main text), where for the checker board model $f_{0}=30$ (e) and for the regular lattice $f_{0}=10.9$ (f).
+(g) $R_{\text{dc}}$ for $f/f^{\prime}=10.9$ (no Fraunhofer) for decreasing bias current.
+Data shifted for clarification.
+(h) Equilibrium loop current configuration ($I=0$) for different values of $f,f^{\prime}$.
+The dots indicate counter- (red) or clockwise (blue) going currents and their size represents the magnitude of the loop current (relative linear scale).
+We now include the Fraunhofer effect, $I_{c}\sim\text{sinc}(\pi f/f_{0})$, where the parameter $f_{0}$ captures the junction area.
+The resulting reduction of the critical current at finite $f$ leads to a base offset in the $R_{\text{dc}}$-curve, see Fig.
+3 e) [the red curve represents Eq.
+(1) with $f$-dependent $I_{c}$], in remarkable resemblance to the experimental data, Fig.
+2 c).
+As a comparison, take the regular square lattice, $f=f^{\prime}$, where a qualitatively very similar beating pattern arises due to Fraunhofer [Fig.
+3 d) versus Fig.
+3 f)], by setting $f_{0}$ (instead of $f/f^{\prime}$) to $\approx 10.9$.
+However, this alternative model can be safely excluded to explain the experimentally observed phenomenology.
+First, the Fraunhofer beating pattern in Fig.
+3 f) skips a beat at $f=0$, due to the sinc-function having no zero at the origin, such that there are two beating pattern frequencies, $2/f_{0}$ (main beat around $f=0$) and $1/f_{0}$ (side beats) [see Fig.
+3 f)].
+Figure 2 only exhibits one beating frequency, in alignment with the checker board model 3 c).
+Moreover, the area ratio extracted from micrographs of the device align much better with the first model ($f/f^{\prime}\approx 10.9$ and much larger $\sim f_{0}$).
+Related to that, for $f/f_{0}\ll 1$ (while $I\gg JI_{c}$), the base offset is approximately linear, $R_{\text{dc}}\sim f$, consistent with the measured asymptotics in Fig.
+2.
+Indeed, nodes separating two beats [green dots in Figs.
+3 e) and f)] can only exhibit an increase in $R_{\text{dc}}$ in the checker board model, while in the regular lattice, the onset of the beating pattern must coincide with $R_{\text{dc}}$ reaching the constant plateau value $\approx R$.
+We observe that the pattern in Fig.
+3 a) transitions from resistance dips to peaks at integer $f+f^{\prime}$, depending on whether $f$ and $f^{\prime}$ are individually closer to integer or half-integer.
+Those peaks can become dips when lowering the bias current, see Fig.
+3 g), marking the onset of the checker board version of regular frustration – which goes hand in hand with the formation of additional resistance dips at $f+f^{\prime}$ half integer [see, $10<f<$ 1 in Fig.
+3 g)]  [^3] Such extra features have not been detected in experiment, likely due to application of higher bias currents.
+Future experiments could be dedicated to the search of additional frustration features..
+Overall, the ’frustrated frustration’ pattern ($f+f^{\prime}\in\mathbb{Z}$) is more stable than regular frustration.
+For an intuitive understanding, consider the equilibrium ($I=0$) configurations of loop currents  [^4] Our definition of loop currents, as shown in Fig.
+3 h), is related to, but distinct from, the more commonly considered vortex picture.
+flowing through single plaquettes, see Fig.
+3 h).
+Thus we find that, e.g., the resistance dip at $f\approx 2/3,f^{\prime}\approx 1/3$ resembles much more closely the configuration of $f=f^{\prime}=1/2$ than the one at $f=f^{\prime}=1/3$ – however, with an overall reduced loop current magnitude, indicating increased stability.
+Stability considerations of the various frustration features can also be understood in terms of the regular vortex model [55].
+Conclusion.
+We demonstrated an array made with 4TJJs.
+For this we introduced an in-situ fabrication technique for arrays.
+A frustration pattern was measured which differs from the expectation for ordinary 2TJJ arrays.
+The difference could be explained by theoretically introducing a checker board lattice with two alternating flux patterns $f$ and $f^{\prime}$.
+The periodicity of the beating pattern can be connected to the area ratios $A$/$A^{\prime}$.
+In particular, while dips in the dc resistance can be linked to the sum $f+f^{\prime}$ being integer, the in general irrational area ratio leads to an irrational repartition of this total flux into the alternating plaquettes.
+This allows for the stabilization of the superconducting BKT phase at irrational fractions of flux penetrating individual plaquettes.
+This feature allows us to estimate the spatial expansion of the 4TJJ central weak link region, the result of which is compatible with the junction geometry.
+Overall, the here considered setup opens up a previously unexplored class of array system with alternating flux textures.
+Given that detailed understanding of vortex lattice structures and their dynamics are still an open question for regular lattices with a single flux $f$ [2], the here introduced alternating $f,f^{\prime}$ lattice will likely open up a new research direction in its own right (e.g., exploring the impact of capacitive quantum fluctuations on the phase diagram as in Ref.
+[8], now including incommensurate vortex features).
+Acknowledgements.
+We thank Herbert Kertz for technical assistance.
+All samples have been prepared at the Helmholtz Nano Facility [60].
+This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769, by the German Federal Ministry of Education and Research (BMBF) via the Quantum Futur project ‘MajoranaChips’ (Grant No.
+13N15264) and the Quantum Futur project with grant no.
+13N14891 (RR), as well as the Bavarian Ministry of Economic Affairs, Regional Development and Energy within Bavaria’s High-Tech Agenda Project ”Bausteine für das Quantencomputing auf Basis topologischer Materialien mit experimentellen und theoretischen Ansätzen” (grant no.
+07 02/686 58/1/21 1/22 2/23).
+SUPPLEMENTAL NOTE 1: Fabrication procedure
+The stencil lithography fabrication process of the four-terminal Josephson junction (4TJJ) array with molecular beam epitaxy is described in the following.
 As a first step, a Si(111) substrate is covered by a 300-nm-thick SiO2 layer followed by a 100-nm-thick Si3N4 layer.
@@
 As a next step, the stencil mask is under-etched by hydrofluoric acid, removing the SiO2 layer below and revealing the Si(111) substrate surface.
-The stencil mask can be seen in Fig.
-1 d).
+The stencil mask can be seen in main text Fig.
+1 c).
 After inserting the pre-patterned substrate into an ultra high vacuum growth chamber, a 10-nm-thick Al2O3 interlayer is first deposited under rotation.
@@
 As a last step, a 5-nm-thick Al2O3 capping layer is deposited under rotation.
-In Fig.
-1 d) a scanning electron microscope image of the final device structure is shown.
+In main text Fig.
+1 c), a scanning electron microscope image of the final device structure is shown.
 The four-terminal Josephson junction consists of four superconducting Nb islands connected at their tips by a central Pt weak link area.
 After the material deposition, the stencil mask is removed by applying ultrasonic power.
-The square unit cell of the array formed by a 4TJJ in each corner is indicated in Fig.
-1 d).
-To prevent screening of magnetic fluxes by screening currents in superconducting material surrounding the array, the Nb layer outside the device is removed by reactive ion etching.
-Figure 1 e) shows a scanning electron microscope image of the 30$\times$30 4TJJ array investigated in this study.
-In addition to the device presented in the main text, we measured an identical array device and a reference two-terminal Josephson junction, all fabricated on the same substrate during the same fabrication run, presented in the supplementary material.
+The square unit cell of the array formed by a 4TJJ in each corner is indicated in main text Fig.
+1 c) as well.
+To prevent screening of magnetic fluxes by screening currents in superconducting material surrounding the array, the Nb layer outside the device is removed by reactive ion etching (see suppl.
+section 7).
+SUPPLEMENTAL NOTE 2: Measurement setup
 The measurements have been performed in a dilution refrigerator with a base temperature of 70 mK equipped with a superconducting magnet.
@@
 The voltage signal, from which the resistance ($R_{\text{dc}}$) and differential resistance ($dV/dI$) has been calculated, was measured with a quasi 4-point setup by applying a dc bias to the array.
-In additional measurements, the dc signal was superimposed with a smaller AC signal in order to measure the differential resistance using the lock-in technique (see supplementary material).
-Magnetotransport measurements have been performed on a $30\times 30$ 4TJJ array at a temperature of 80 mK.
-At zero magnetic field, the array shows a critical current of $I_{\mathrm{c}}=57\;\mu$A and, close to the superconducting regime, the device has a differential resistance of around 5.5 $\Omega$ (see supplementary material).
-The differential resistance-bias current-magnetic field diagram depicted in Fig.
-2 a) shows periodic oscillations of the critical current with magnetic field.
-When applying a direct current of 30 $\mu$A through the device, its resistance oscillates with the same periodicity in magnetic field (cf.
-Fig.
-2 b), which was determined to be 6.25 mT by a fast Fourier transform (see supplementary material).
-In addition, a device- and array-independent (present in the reference 2TJJ and both arrays) magnetic hysteresis of the resistance pattern was measured, which is not further considered in the main text (see supplementary material).
-[…]
-The resistance signal is similar to what, e.g.
-Rzchowski et al.
-[2] measured, however, the expected frustration pattern for rational values of $f$ is missing.
-As can be seen in Fig.
-2 c), at certain magnetic field around ±$\pm 5\;\Phi_{0}/A_{\mathrm{uc}}$, the oscillations first disappear and then reappear when further increasing the magnetic field.
-As will be explained below, both features, the different frustration pattern and the beating pattern, stem from the altered array lattice, which can be described by a conventional 2TJJ square lattice with checker board flux pattern.
-This gives rise to a superimposed beating pattern in addition to the flux quantum oscillations of the array unit cell.
-The periodicity of that beating pattern is approximately equal to the area ratio of the large array unit cell ($A$) and the 4TJJ central weak link region ($A^{\prime}$).
-The base offset of the resistance dips can be explained by the Fraunhofer effect of the 2TJJs between two individual contacts of a 4TJJ.
-The resistance oscillations can be seen to magnetic fields above ±$\pm 100\;$mT with, in total, 30 flux quantum oscillations ($30\;\Phi_{0}/A_{\mathrm{uc}}$).
-Figure 2: Measurement data of the 30$\times$30 4TJJ array at 80 mK.
-a) Differential resistance as a function of bias current and magnetic field.
-A periodic oscillation of the critical current is clearly visible.
-b) Resistance as a function of magnetic field with an applied dc bias of 30 $\mu$A.
-The resistance oscillations correspond to a periodicity of 6.25 mT.
-At around ±$\pm 5\;\Phi_{0}/A_{\mathrm{uc}}$, the resistance oscillations are damped.
-c) Resistance oscillations under magnetic field up to ±$\pm 140$ mT with an applied dc bias of 30 $\mu$A.
-For magnetic fields higher than the point of damping, the oscillations stay present to above $\pm$100 mT.
-To explain the measured behavior of the 4TJJ array, a corresponding theoretical model is introduced in the following.
-The experimental device is schematically depicted in Fig.
-1 a).
-Its square unit cell has one 4TJJ in each corner.
-The junctions are connected via their superconducting Nb contacts and every junction has a central Pt weak link region.
-Based on the RCSJ network model for multi-terminal Josephson junctions [28], the array can be described as shown in Fig.
-1 b) with four 2TJJs creating one 4TJJ (see also supplementary material).
-This introduces a second lattice of areas pierced by magnetic flux, the central weak link region of the 4TJJs.
-The multi-terminal array model as shown schematically in Figs.
-1 a) and b) can be conveniently represented in terms of an ordinary square lattice junction model [see Fig.
-1 c)], with the following important difference to previous theoretical and experimental studies: instead of all the plaquettes being pierced by the same flux, we get an alternating (checker board) flux pattern ($f=\Phi/\Phi_{0}$ and $f^{\prime}=\Phi^{\prime}/\Phi_{0}$).
-We capture the dynamics of this system by means of a classical RCSJ model approach, where the equations of motion for the superconducting phases $\phi_{j}$ at nodes $j=(j_{x},j_{y})$ follow as usual from the Kirchhoff laws (for more details see supplementary material).
-Including a source and drain contact on two sides of the lattice, the RCSJ model allows for a direct computation of the dc resistance $R_{\text{dc}}$ as a function of the magnetic field.
-We do so by an explicit numerical evaluation of the classical RCSJ equations of motion.
-In addition we track the state of the system by a quantity related to the more commonly considered vorticity [4]: for a given plaquette, we calculate the counter clockwise supercurrent in the loop defining the plaquette (where for two neighbouring circuit nodes $\phi_{i},\phi_{j}$, the supercurrent is given as $I_{c}\sin(\phi_{i}-\phi_{j}+a_{ij})$ with the equivalent of the Peierl’s phase $a_{ij}$ capturing the magnetic flux), see also Fig.
-3 a).
-This loop current captures the vortex structure qualitatively quite well, see, e.g., the loop current texture of the regular square lattice system $f^{\prime}=f$ for $f=1/2,1/3$ in Figs.
-3 b) and c).
-Contrary to the vorticity however, the loop current does not return an integer number.
-This property will turn out to be useful in a moment, to describe stable states for the checker board flux system, $f^{\prime}\neq f$.
-Figure 3: Results of the theoretical analysis.
-a) Checker board arrangement of alternating fluxes $f,f^{\prime}$, and depiction of superconducting loop currents (inset within green circle) as an alternative quantity to represent the vortex configuration.
-While loop currents do not provide an integer-valued quantity (contrary to vorticity), the ordinary $f=1/2$ (b) and $f=1/3$ (c) states are nontheless clearly visible by structures repeating either over $2$ or $3$ plaquettes, respectively.
-The dots indicate clockwise (blue) or counterclockwise (red) going currents and their size represents the magnitude of the loop current (relative linear scale).
-For the checker board ($f\neq f^{\prime}$), special points are those where $f+f^{\prime}$ adds up to an integer (d).
-At these points, the dc resistance (e) experiences superconducting dips and exhibits a beating pattern with large periodicity, consistent with experimental data.
-The current loop texture for the stationary states (no bias current) at the respective dips is shown in panels (f-i), which differ markedly from the usual stable (or semistable) states of conventional lattices with $f=f^{\prime}$ (b,c).
-That being said, we explicitly compute $R_{\text{dc}}$ as a function of $f$ for a square lattice with $6\times 6$ nodes  [^1] The theory modelling occurs with a smaller system compared to the actual experimental setup.
-We do so simply to reduce the computational effort; the qualitative features of the theory model do not depend strongly on system size.
-for $f/f^{\prime}=5.4$, yielding Fig.
-3 e).
-The central finding is this: the dc resistance experiences superconducting dips whenever $f=-f^{\prime}$ modulus $1$, see Fig.
-3 d).
-Consequently, while the flux through a single plaquette is far from being zero at a given dip (or an integer of the flux quantum), the sum of two neighbouring plaquettes $f+f^{\prime}$ is.
-Denoting the ratio $f^{\prime}/f=\beta$, these dips occur for $f_{m}=m/(1+\beta)$ with $m\in\mathbb{Z}$.
-Since $\beta$ is given by the area ratio (see discussion of the experimental section above), $\beta$ is in general an irrational number, such that the dips in $R_{\text{dc}}$ likewise occur at irrational values, $f_{m}\notin\mathbb{Q}$.
-This is in stark contrast to regular square lattices commonly considered in previous works ($\beta=1$), where $R_{\text{dc}}$ dips at integer $f$, with weaker side dips for rational $f$ with low denominator.
-In addition to the dips repeating with period $p_{\text{dips}}=1/(\beta+1)$ in $f$ space, there is an overall beating pattern due to the fact that there repeatedly occur configurations where both $f_{m}$ and $\beta f_{m}$ (value of $f^{\prime}$ at the dip) return to a value close to zero (up to modulus $1$), where we approximately recover the state at zero applied field.
-The period for that pattern is $p_{\text{beating}}\approx 1/\beta$, where, due to $\beta\ll 1$ (large discrepancy between two relevant areas in experimental setup) $p_{\text{beating}}\gg p_{\text{dip}}$.
-This finding agrees very well with the experimental data shown in Figs.
-2 b) and c).
-When converting the frustration dip periodicity into magnetic field dip periodicity $p_{\text{dips,B}}$ (6.25 mT) with $1/(\beta+1)=p_{\text{dips,B}}\cdot A/\Phi_{0}$ and simplifying, we find $\Phi_{0}/p_{\text{dips,B}}=A^{\prime}+A=A_{\text{uc}}$.
-This result agrees with Fig.
-2 b).
-In case that $A^{\prime}$ and $A$ are unknown, one can use the the difference in periodicities $p_{\text{beating}}-p_{\text{dips}}\approx 1/\beta-1/(\beta+1)$ to estimate $\beta$.
-Here, a value of $\beta\approx 0.092$ was determined (see supplementary material).
-The area of the central weak link region is given by $A^{\prime}=A\beta=(A_{\text{uc}}-A^{\prime})\beta$, resulting in
-$A^{\prime}=A_{\text{uc}}\frac{\beta}{1+\beta}.$ (1)
-With this, the central weak link region of the 4TJJs is $A^{\prime}\approx(165\;\text{nm})^{2}$. This is compatible with the 4TJJ geometry (see supplementary material).
-The only feature that is not captured is the base offset of the resistance dips that occurs for some finite magnetic fields (the fact that the $R_{\text{dc}}$ dips return to a finite value instead of zero).
-This extra phenomenology can however be explained very easily: the above deployed RCSJ model assumes critical currents for the individual junctions that are independent of the magnetic field.
-In reality, given the estimated areas of the junctions, there is a finite onset of a Fraunhofer pattern, reducing the critical current at stronger magnetic fields (Fig.
-2 a).
-When biasing the array with a constant current and ramping the magnetic field, this results in the finite-value- instead of zero-resistance dips.
-Ultimately, the checker board arrangement of alternating fluxes ($f,f^{\prime}$) provides a unique opportunity to stabilize the superconducting (insulating) phase of the Berezinskii–Kosterlitz–Thouless (BKT) transition for incommensurate (irrational) flux quanta, since two neighbouring plaquettes can be tuned to values that cancel (up to modulus $1$) over a unit cell containing two plaquettes.
-The insulating phase of the vortex BKT transition corresponds to a superconducting phase in terms of the transport properties (which is why we refer to this phase as the superconducting phase for the remainder of this work).
-The special feature of stabilizing the superconducting phase for incommensurate (irrational) flux quanta motivates us to study the stationary system state (at zero bias current) at the values $f_{m}$ where $R_{\text{dc}}=0$ (suppressed vortex motion).
-Due to the irrational values for $f,f^{\prime}$, the (integer-valued) vorticity is not well-suited for such an analysis.
-This is why we instead compute the steady state configuration of the previously introduced loop currents in Figs.
-3 f)-i).
-Even though $f,f^{\prime}\notin\mathbb{Q}$, the unit cell of the stationary state is always of the size of two plaquettes, due to $f+f^{\prime}$ cancelling – in contrast to, e.g., the regular $f=f^{\prime}=1/3$ state, which repeats over three plaquettes.
-In fact, the loop current texture of the respective stationary states gradually changes between the regular $f=f^{\prime}=0$ state (when both $f$ and $f^{\prime}$ are roughly integer) and the regular $f=f^{\prime}=1/2$ state (note that for half-integer flux quanta, $1/2$ is equivalent to $-1/2$, hence the checker board state with $f=-f^{\prime}$ is here indistinguishable from $f=+f^{\prime}$).
-The here presented work thus provides a glimpse into classes of states that have not been investigated so far.
-While the dips at the beginning and the end of a beating pattern (period $p_{\text{beating}}$) correspond very closely to the regular states $f=f^{\prime}$ integer, the dip in the middle of the beating period corresponds to $f=f^{\prime}=1/2$ – thus evidencing a regular frustration pattern.
-The dips in between provide states that defy the usual frustration picture as they occur at flux values far away from rational flux quanta.
-Therefore, we can think of the checker board arrangement with incommensurate ratio $\beta$ as a way to circumvent usual frustration.
-Finally, the above finding can be cast into the more commonly used language of vortices and the BKT transition.
-The energy of the classical 2D Josephson junction array in terms of the number of vortices $v$ per plaquette is given by (see, e.g., Ref.
-[9] for details),
-$H=\sum_{i,j}(v_{i}-f_{i})U_{ij}(v_{j}-f_{j})\ ,$ (2)
-where the indices $i,j$ address the plaquette positions in 2 dimensions (i.e., both $i$ and $j$ implicitly encode a pair of $x$ and $y$ coordinates), and $U_{ij}$ represents the typical logarithmic long-range interaction.
-For regular square lattices $f_{i}=f$ (i.e., all plaquettes experience the same flux), the well-known BKT picture arises.
-At zero flux $f=0$, the zero vortex state $v_{j}=0$ (which is insulating in that no vortices are available to move) corresponds to the ground state, whereas states with finite vortex numbers or vortex pairs (conducting states) are at low temperatures inaccessible due to a finite energy gap, leading to the BKT transition.
-At finite flux, e.g., $f>0$, the system reacts in such a way that the new ground state can be to a good approximation described by a state with a finite vortex density of $\sim f$ vortices per plaquette, in order to screen the applied external flux, where (as is well-known) the propensity of the ground state to have ordered (and thus likewise insulating) vortex configurations or many freely moving vortex configurations depends on whether $f$ is a rational number or not.
-For the here present checker board flux texture ($f_{i}=f,f^{\prime}$ for alternating plaquette positions) the situation is different in the following crucial aspect.
-The aforementioned special points of cancelling neighbouring fluxes $f^{\prime}=-f$ still have the zero vortex state $v_{j}=0$ as the gapped ground state because averaged over the whole system, no flux is applied, and hence no finite vortex density occurs due to screening.
-We therefore arrive again at the superconducting BKT phase, notably, at generally irrational values of $f$.
-We emphasise that these zero flux states are by no means trivial.
-Consider the special (rational) case of $f=1/2$ and $f^{\prime}=-1/2$, with the zero vortex ground state.
-With a simple unitary transformation (increasing the vortex number on every $-1/2$ plaquette by one), we map to the regular square lattice $f=f^{\prime}=1/2$ with the well-known ground state given by an alternating $0,1$ vortex configuration.
-Consequently, $f=-f^{\prime}=1/2$ with the zero vortex ground state and $f=f^{\prime}=1/2$ with alternating $0,1$ vortex ground state are equivalent.
-Now, the difference to the case where $f=-f^{\prime}$ but $f$ irrational (or in fact any rational or irrational number other than $f=0$ or $f=1/2$) is that the correspondence to the regular lattice $f=f^{\prime}$ is no longer exact (as there exists no unitary mapping due to $v_{i}\in\mathbb{Z}$).
-For values close to $1/2$, there still exists a unitary map to a system approximately equivalent to $f=f^{\prime}=1/2$, this correspondence grows weaker and weaker as we move away from $1/2$, which at the value of $1/4$ or lower leads to a cross over, where the checker board system with $f=-f^{\prime}$ more closely imitates the regular $f=f^{\prime}=0$ system instead of $f=f^{\prime}=1/2$.
-This constitutes a vortex perspective which is ultimately equivalent to the above analysis in terms of superconducting loop current configurations changing from the $f=f^{\prime}=0$ to the $f=f^{\prime}=1/2$ ground state [see also Figs.
-3 f)-i) and their discussion above].
-We demonstrated an array made with 4TJJs.
-For this we introduced an in-situ fabrication technique for array fabrication.
-A frustration pattern was measured which differs from the expectation for ordinary 2TJJ arrays.
-The difference could be explained by theoretically introducing a checker board lattice with two alternating flux patterns $f$ and $f^{\prime}$.
-The periodicity of the beating pattern can be connected to the area ratios $A$/$A^{\prime}$.
-In particular, while dips in the dc resistance can be linked to the sum $f+f^{\prime}$ being integer, the in general irrational area ratio leads to an irrational repartition of this total flux into the alternating plaquettes.
-This allows for the stabilization of the superconducting BKT phase at irrational fractions of flux penetrating individual plaquettes – a feature we termed frustrated frustration.
-This feature allows us to estimate the spatial expansion of the 4TJJ central weak link region, the result of which is compatible with the junction geometry.
-Overall, the here considered setup opens up a previously unexplored class of array system with alternating flux textures.
-Given that detailed understanding of vortex lattice structures and their dynamics are still an open question for regular lattices with a single flux $f$ [9], the here introduced alternating $f,f^{\prime}$ lattice will likely lead to a rich catalogue of open research directions of its own, the full extent of which we could not possibly anticipate at this stage.
-As an example of a practical near-term question, we expect that at lower critical currents it might be perceivable to measure signatures related to rational (instead of integer) values of the sum $f+f^{\prime}$, leading to substructures within the beating pattern profile.
-Also, non-local (e.g., multi-terminal) transport on such arrays, as well as their Shapiro response, could reveal further information about any nontrivial vortex configuration or motion.
-We thank Herbert Kertz for technical assistance.
-All samples have been prepared at the Helmholtz Nano Facility [33].
-This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769, by the German Federal Ministry of Education and Research (BMBF) via the Quantum Futur project ‘MajoranaChips’ (Grant No.
-13N15264), as well as the Bavarian Ministry of Economic Affairs, Regional Development and Energy within Bavaria’s High-Tech Agenda Project ”Bausteine für das Quantencomputing auf Basis topologischer Materialien mit experimentellen und theoretischen Ansätzen” (grant no.
-07 02/686 58/1/21 1/22 2/23).
-SUPPLEMENTAL NOTE 1: Second 30$\times$30 4TJJ Array
+In additional measurements (suppl.
+section 9), the dc signal was superimposed with a smaller AC signal in order to measure the differential resistance using the lock-in technique and the ac signal has been used alone for ac resistance measurements dependent on magnetic field.
+SUPPLEMENTAL NOTE 3: Second 30$\times$30 4TJJ array
 In addition to the 30$\times$30 4TJJ array presented in the main text, there was a second device on the same sample.
@@
 The critical current oscillates with magnetic field in the same way as for the main text array.
-With a DC bias of 30 $\mu$A, that behavior leads to an oscillation of the resistance with magnetic field (see Fig.
-S4 b).
+With a DC bias of 30 $\mu$A, that behavior leads to an oscillation of the resistance with magnetic field [see Fig.
+S4 a)].
 Similar to the main text device, the frustrated frustration pattern is present.
 The damping of the oscillations of resistance and critical current can be reproduced.
-At zero magnetic field and a temperature of 80 mK, the critical current is comparable to the main text device with $I_{\mathrm{c}}=59.75\;\mu$A (see Fig.
-S1 c and d).
+At zero magnetic field and a temperature of 80 mK, the critical current is comparable to the main text device with $I_{\mathrm{c}}=59.75\;\mu$A [see Fig.
+S1 c) and d)].
 The differential resistance is lower with $R=1.75\;\Omega$ close to the maximum applied current of 100 $\mu$A.
 Supplemental Figure S1: Data of the second 30$\times$30 four-terminal Josephson junction array.
-a) Scanning electron microscopy image of the device.
-b) Differential resistance-bias current-magnetic field diagram.
+a) Differential resistance-bias current-magnetic field diagram.
 The oscillations of the critical current are clearly visible.
+b) Scanning electron microscopy image of the device.
 c) I-V measurement at 80 mK.
-d) Differential resistance calculated from the data in figure c).
-SUPPLEMENTAL NOTE 2: Reference two-terminal Josephson junction
-On the chip of the presented arrays, a reference two-terminal Josephson junction has been fabricated in the same way as the arrays (see figure S2).
+d) Differential resistance calculated from the data in c).
+SUPPLEMENTAL NOTE 4: Reference two-terminal Josephson junction
+On the chip with the presented arrays, a reference two-terminal Josephson junction has been fabricated in the same way as the arrays (see Fig.
+S2).
 The weak link length was determined to be 92 nm and its width is 920 nm.
-Although only a width of 809 nm is indicated in figure S2, Pt and Nb material under the edge of the bottom Si3N4 layer needs to be taken into account.
-The angle between the substrate surface and niobium beam was 57.7.
+Although only a width of 809 nm is indicated in Fig.
+S2, Pt and Nb material under the edge of the bottom Si3N4 layer needs to be taken into account.
+The angle between the substrate surface and Nb beam was 57.7.
 The 300 nm-thick-SiO2 layer right at the edge of the trench is etched away, just like under the mask itself.
 This leaves the 100 nm-thick-Si3N4 layer under-etched at the edge and gives additional substrate space for the niobium shadow to be deposited.
-The extent of the Nb shadow under the Si3N4 edge can be derived from the top shadow of image S2 (174 nm).
+The extent of the Nb shadow under the Si3N4 edge can be derived from the top shadow of Fig.
+S2 (174 nm).
 The bottom shadow has a smaller distance from its edge than the top shadow because it is delimited by the lower edge of the Si3N4 layer (instead of the upper edge).
@@
 This results in a total bottom shadow of 111 nm, which has already been added to the width specification of this Josephson junction above (920 nm).
-In figure S2, the junction is shifted with respect to the stencil mask (by 42.2) because latter was rotated against the niobium beam during deposition.
+In Fig.
+S2, the junction is shifted horizontally with respect to the stencil mask [by $174\;\text{nm}\cdot\text{tan}(42.2^{\circ})=158\;\text{nm}$] because latter was rotated against the Nb beam during deposition.
 Supplemental Figure S2: Scanning electron microscope images of the reference two-terminal Josephson junction.
 Figure S3 shows the voltage and differential resistance dependent on the current.
-The device has a critical current of 56.51 $\mu$A and a differential resistance of 1.21 $\Omega$.
+The device has a critical current of $I_{\text{c}}=56.51\;\mu$A and a differential resistance of $R=1.21\;\Omega$.
 The Josephson energy of the junction is $E_{\mathrm{J}}=\Phi_{0}I_{\mathrm{c}}/2\pi=116$ meV.
@@
 […]
+The Stewart-McCumber parameter for these values (assuming $R_{N}\approx R$) is $\beta_{\text{c}}=2eI_{\text{c}}R_{\text{N}}^{2}C/\hbar\approx 1.4\cdot 10^{-5}$.
+Therefore, the junction is in the over-damped regime.
+Although the junction is over-damped, the current-voltage characteristic of the device shows hysteretic behavior [see Fig.
+S3 a)].
+This can be explained by an increase in electron temperature in the resistive state [1].
 […]
-SUPPLEMENTAL NOTE 3: Magnetic hysteresis of the resistance pattern
+SUPPLEMENTAL NOTE 5: Magnetic hysteresis of the resistance pattern
 A magnetic hyteresis has been observed when measuring the resistance with magnetic field.
@@
 A possible reason is the screening of magnetic field by present Abrikosov vortices.
+In the main text, we remove the hysteresis by showing only one sweep direction of the magnetic field, namely from negative to positive, and shift the data by 0.9 mT to the negative side.
+The period of the Fraunhofer effect for the reference two-terminal Josephson junction (2TJJ) in Fig.
+S4 c) does not fit to the weak link area given in suppl.
+section 4.
+Possible reasons for this are an unequal current distribution, flux focusing, or the presence of Abrikosov vortices.
 (a) Main text 4TJJ array.
@@
 All devices show hysteretic behavior with magnetic field.
-The two upper resistance oscillations stem from the frustration pattern of the array, the lower resistance oscillation stems from the superconducting quantum interference of the two-terminal Josephson junction.
-SUPPLEMENTAL NOTE 4: Determination of the resistance oscillation period
+The two upper resistance oscillations stem from the frustration pattern of the array, the lower resistance oscillation stems from the Fraunhofer effect of the two-terminal Josephson junction.
+SUPPLEMENTAL NOTE 6: Determination of the resistance oscillation period
 For the determination of the oscillation frequencies, fast Fourier transforms have been applied to the data in Fig.
-3 c) and d) of the main text (see Fig.
+2 b) and c) of the main text (see Fig.
 S5).
 The frequency spectra of both data curves match.
-The dominant frequency (period) is 0.16 1/mT (6.25 mT).
+The dominant frequency (period) is 0.16 mT-1 (6.25 mT).
 The period of the beating pattern cannot be seen in the spectrum, probably because it is too close to the zero peak.
 Supplemental Figure S5: Fast Fourier transforms of the data curves in main text Fig.
-3 c) (fine sweep) and d) (wide range).
-The dominant frequency (period) is 0.16 1/mT (6.25 mT).
+2 b) (fine sweep) and c) (wide range).
+The dominant magnetic field frequency (period) is 0.16 mT-1 (6.25 mT).
 The other peaks of the spectrum are higher harmonics of that frequency.
-SUPPLEMENTAL NOTE 5: Description of the array device
+SUPPLEMENTAL NOTE 7: Description of the array device
 This section gives a detailed description of the array device in the main text, including geometrical properties of the array lattice (see Fig.
 S6).
-According to figure S6 a), the unit cell of the array is a square of size $A_{\text{UC}}=573\;\text{nm}\cdot 567\;\text{nm}=(570\;\text{nm})^{2}$.
-The geometry of the weak link area between two neighboring leads of one 4TJJ (see Fig.
-S6 b) is described, on average, by the following dimensions: The width of the weak link is 129 nm and its length is 97 nm.
-The 4TJJ leads have a length of 471 nm.
+According to Fig.
+S6 a), the unit cell of the array is a square of size $A_{\text{UC}}=573\;\text{nm}\cdot 567\;\text{nm}=(570\;\text{nm})^{2}$.
+The geometry of the weak link area between two neighboring leads of one 4TJJ [see Fig.
+S6 b)] is described, on average, by the following dimensions: The width of the weak link is 129 nm and its length is 97 nm.
+The superconducting arms in the 4TJJ array have a length of 471 nm.
 During fabrication of the device, a trench is etched into the SiO2 (300 nm) / Si2N3 (100 nm) layer, down to the Si(111) substrate.
@@
 To minimize disturbance effects, the superconductor around the device is etched away as a last fabrication step, leaving an edge of superconductor of approximately 1 µm around the trench behind.
-To prevent this closed loop of superconductor from generating screening currents, the edge is disconnected close to the ends of the device contacts (see figure S6 c-d).
+To prevent this closed loop of superconductor from generating screening currents, the edge is disconnected close to the ends of the device contacts [see Figs.
+S6 c) and d)].
 Supplemental Figure S6: Scanning electron microscope images of the main text array.
@@
 The left side is the SiO2 (300 nm) / Si2N3 (100 nm) layer around the device.
-SUPPLEMENTAL NOTE 6: Characteristic RCSJ values between two leads of a 4TJJ in the arrray
+SUPPLEMENTAL NOTE 8: Characteristic resistively capacitively shunted junction (RCSJ) values between two leads of a 4TJJ in the arrray
 In this section, relevant parameters for the theoretical description of the array are derived from geometrical considerations of the prior section, as well as from measurement data of the reference two-terminal Josephson junction.
-We assume that critical current density and resistivity, derived in section 2, are geometry independent.
-We then use these values, together with the 4TJJ geometry to derive capacitance, resistivity, and critical current between two leads of one 4TJJ.
-The coplanar capacitance is calculated using the geometrical values given above.
+We assume that critical current density and resistivity, derived in suppl.
+section 4, are geometry independent.
+We then use these values, together with the 4TJJ geometry, to derive capacitance, resistance, and critical current between two leads of one 4TJJ.
+The coplanar capacitance is calculated using the geometrical values given in suppl.
+section 7.
 The relative permeability $\epsilon_{r}$ is assumed to be one, since the weak link is a metal.
@@
 The resistance is equal to $R=9.1\;\Omega$.
-SUPPLEMENTAL NOTE 7: Additional measurements of the main text 30$\times$30 4TJJ array
+The Stewart-McCumber parameter for these values (assuming $R_{N}\approx R$) is $\beta_{\text{c}}=2eI_{\text{c}}R_{\text{N}}^{2}C/\hbar\approx 5.4\cdot 10^{-6}$.
+Therefore, the junction is in the over-damped regime.
+Although the junction is over-damped, the current-voltage characteristic of the array shows hysteretic behavior [see Fig.
+S7 a)].
+This can be explained by an increase in electron temperature in the resistive state [1].
+SUPPLEMENTAL NOTE 9: Additional measurements of the main text 30$\times$30 4TJJ array
 The current-voltage characteristics of the device presented in the main text can be seen in Fig.
@@
 The device has a critical current of $57\;\mu$A and a differential resistance of around 5.5 $\Omega$ close to the maximum applied current of 100 $\mu$A.
-Supplemental Figure S7: IV measurements at 80 mK.
+Supplemental Figure S7: IV measurements of the array discussed in the main text at 80 mK.
 a) The DC IV curve with a critical current of $I_{\mathrm{c}}=57\;\mu$A.
@@
 The resulting resistance is lower than with the DC measurement because the time-dependent sinusoidal signal includes the superconducting state of the device as well.
-When comparing figure S9 a) with figure S4 a), one can see that the oscillations are identical in magnetic field, although the peak pattern is more symmetrical with the AC method.
+When comparing Fig.
+S9 a) with Fig.
+S4 a), one can see that the oscillations are identical in magnetic field, although the peak pattern is more symmetrical with the AC method.
 (a) $I_{\mathrm{AC}}=30\;\mu$A.
@@
 The blue lines indicate a magnetic sweep direction from negative to positive and the orange lines from positive to negative.
-Figure LABEL:fig:SUPP_MISC_RB_DC_I_T a) shows resistance measurements dependent on magnetic field and DC current.
+Figure S10 a) shows resistance measurements dependent on magnetic field and DC current.
 At higher currents, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
-Figure LABEL:fig:SUPP_MISC_RB_DC_I_T b) shows resistance measurements dependent on magnetic field and temperature.
+Figure S10 b) shows resistance measurements dependent on magnetic field and temperature.
 At higher temperatures, superconductivity is not realized at any magnetic field and the oscillation amplitudes decline.
@@
 b): At a DC bias of $I_{\mathrm{DC}}=30\;\mu$A and temperatures between 70 mK and 800 mK.
-SUPPLEMENTAL NOTE 8: Details of the theoretical modelling
+SUPPLEMENTAL NOTE 10: Details of the resistively capacitively shunted junction model calculations
 We here summarize the model calculations for the theoretical discussion of the main text.
-We describe the multiterminal Josephson junction array in the overdamped regime (discarding capacitive contributions), we get at each node $j$ the equation of motion,
-$\frac{1}{2eR}\sum_{\langle j^{\prime}\rangle_{j}}\left(\dot{\phi}_{j}-\dot{\phi}_{j^{\prime}}\right)=-I_{c}\sum_{\langle j^{\prime}\rangle_{j}}\sin(\phi_{j}-\phi_{j^{\prime}}+a_{jj^{\prime}})+\sum_{\langle j^{\prime}\rangle_{j}}\eta_{jj^{\prime}}(t)\ ,$ (S1)
-with shunt resistance $R$, critical current $I_{c}$, and the current noise term $\eta_{jj^{\prime}}$ due to thermal fluctuations from the resistor connecting nodes $j$ and $j^{\prime}$, where
-$\langle\eta_{ii^{\prime}}(t)\eta_{jj^{\prime}}(t^{\prime})\rangle=\frac{2k_{B}T}{R}\delta_{ij}\delta_{i^{\prime}j^{\prime}}\delta(t-t^{\prime})\ .$ (S2)
+We describe the multiterminal Josephson junction array in the overdamped regime (discarding capacitive contributions), we get at each node $j=(j_{x},j_{y})$ (where $j_{x,y}=1,\ldots,J_{x,y}$) the equation of motion,
+$\frac{1}{2eR}\sum_{\langle j^{\prime}\rangle_{j}}\left(\dot{\phi}_{j}-\dot{\phi}_{j^{\prime}}\right)=-I_{c}\sum_{\langle j^{\prime}\rangle_{j}}\sin(\phi_{j}-\phi_{j^{\prime}}+a_{j,j^{\prime}})+\sum_{\langle j^{\prime}\rangle_{j}}\eta_{j,j^{\prime}}(t)\ ,$ (S1)
+with parallel shunt resistance $R$ and critical current $I_{c}$.
+The vector potential $a_{j,j^{\prime}}$ accounts for the magnetic flux piercing the plaquettes of the array (which will be defined in detail futher below).
+While we assume zero temperature in the device for the discussion in the main text, we here include for completeness the current noise term $\eta_{j,j^{\prime}}$ due to thermal fluctuations from the resistor connecting nodes $j$ and $j^{\prime}$, where
+$\langle\eta_{i,i^{\prime}}(t)\eta_{j,j^{\prime}}(t^{\prime})\rangle=\frac{2k_{B}T}{R}\delta_{i,j}\delta_{i^{\prime},j^{\prime}}\delta(t-t^{\prime})\ .$ (S2)
 The notation $\sum_{\langle j^{\prime}\rangle_{j}}$ indicates that the sum $j^{\prime}$ goes over nearest neighbours of $j$.
-SUPPLEMENTAL NOTE 9: Estimation of the central weak link area
-According to the theoretical discussion in the main text, the parameter $\beta$ and the frustration dip periodicities are defined as:
-$\beta=\frac{f^{\prime}}{f}=\frac{A^{\prime}}{A}$ (S3)
-$p_{\text{dips}}=\frac{1}{\beta+1}$ (S4)
-$p_{\text{beating}}\approx\frac{1}{\beta}$ (S5)
-With this, the difference of the two periodicities can be estimated:
-$\Delta p\approx\frac{1}{\beta}-\frac{1}{\beta+1}$ (S6)
-For this estimate the $\Delta p$ was determined by counting the number of resistance dips between the two beating points in figure 2 c).
-Ten dips ($\Delta p\approx 10$) can be seen between the two beating points.
-This yields $\beta\approx 0.092$ and, thus, an estimated weak link area of $A^{\prime}=A_{\text{UC}}\cdot\beta/(1+\beta)\approx(165\;\text{nm})^{2}$.
-For comparison, such a weak link area is sketched on one 4TJJ in Fig.
+That is, for $j=(j_{x},j_{y})$, its neighbours are $(j_{x}\pm 1,j_{y})$ and $(j_{x},j_{y}\pm 1)$.
+The system further has boundary conditions.
+At the top and bottom boundaries ($j_{y}=1$ and $j_{y}=J_{y}$), the nodes simply terminate, such that the nearest neighbour sum only has three instead of the usual four terms.
+On the left hand side, $j_{x}=1$, the array connects to ground.
+This connection can be taken into account by artifically extending the lower bound of the $j_{x}$ index to 00 (instead of $1$) and setting $\phi_{0,j_{y}}=0$.
+The right-hand side on the other hand ($j_{x}=J_{x}$) connects to the current-biased terminal.
+Here, we artificially extend of the upper bound of $j_{x}$ from $J_{x}$ to $J_{x}+1$ and replace in the sums $\phi_{(J_{x}+1,j_{y})}=\phi_{b}$, where $\phi_{b}$ is the superconducting phase of the current-biased terminal.
+This phase satisfies an equation of motion of its own,
+$\frac{1}{2eR}\sum_{j_{y}}\left(\dot{\phi}_{b}-\dot{\phi}_{(J_{x},j_{y})}\right)=-I_{c}\sum_{j_{y}}\sin\left(\phi_{b}-\phi_{(J_{x},j_{y})}+a_{b,(J_{x},j_{y})}\right)+\sum_{j_{y}}\eta_{b,(J_{x},j_{y})}+I\ ,$ (S3)
+where the current bias $I$ appears as an additional force term.
+The voltage built up across the device is defined through the phase velocity $\dot{\phi}_{b}$.
+The dc resistance across the total device is thus given as
+$R_{\text{dc}}=\lim_{\tau\rightarrow\infty}\frac{\phi_{b}(\tau)-\phi_{b}(0)}{2e\tau I}\ .$ (S4)
+For sufficiently long measurement times $\tau$, this quantity no longer depends on the initial conditions at $t=0$.
+The vector potentials must be chosen such that sums along counter-clockwise loops yield the flux through a plaquette $(j_{x},j_{y})$ (note that plaquette indices are of course shifted with respect to node indices).
+Defining that flux as $f_{(j_{x},j_{y})}$, then the vector potential needs to satisfy
+$f_{(j_{x},j_{y})}=a_{(j_{x},j_{y}),(j_{x}+1,j_{y})}+a_{(j_{x}+1,j_{y}),(j_{x}+1,j_{y}+1)}+a_{(j_{x}+1,j_{y}+1),(j_{x},j_{y}+1)}+a_{(j_{x},j_{y}+1),(j_{x},j_{y})}\ .$ (S5)
+For the checker board model, the flux yields either $f$ or $f^{\prime}$ if $j_{x}+j_{y}$ is even or odd, respectively.
+We further define $a_{j_{x},j_{y}}^{(y)}=a_{(j_{x},j_{y}),(j_{x},j_{y}+1)}$ and $a_{j_{x},j_{y}}^{(x)}=a_{(j_{x},j_{y}),(j_{x}+1,j_{y})}$, and choose $a_{j_{x},j_{y}}^{(y)}=0$.
+Using $a_{jj^{\prime}}=-a_{j^{\prime}j}$, we find a valid gauge as
+$a_{(j_{x},j_{y})}^{(x)}=\sum_{j_{x}^{\prime}=1}^{j_{x}}f_{(j_{x}^{\prime},j_{y})}\ ,$ (S6)
+which corresponds to a lattice version of the Landau gauge.
+In the absence of the noise term, the system can be solved as a first order differential equation, which for all numerical results shown in the main text is done directly with the NDSolve routine on Mathematica.
+In Eq.
+(1) in the main text, we provide an exact expression for the special case of $f,f^{\prime}\in\mathbb{Z}$ (equivalent to $f=f^{\prime}=0$).
+We here derive this expression.
+For the phases within the array, we choose the ansatz
+$\phi_{(j_{x},j_{y})}=\frac{j_{x}}{J_{x}+1}\phi_{b}\ .$ (S7)
+At zero temperature and $f=f^{\prime}=0$, this ansatz immediately solves Eq. (S2) at all nodes. The remaining equation of motion is simply that of $\phi_{b}$, which is simply given as
+$\frac{1}{2eR}\frac{1}{J_{x}+1}\dot{\phi}_{b}=-I_{c}\sin\left(\frac{1}{J_{x}+1}\phi_{b}\right)+\frac{I}{J_{y}}\ .$ (S8)
+This is simply the equation of motion for a single Josephson junction with voltage and bias current rescaled by the lattice size parameters $J_{x}+1,J_{y}$.
+The resulting dc resistance of this reduced equation is well known.
+Carefully accounting for the rescaling, we arrive at Eq.
+(1).
+SUPPLEMENTAL NOTE 11: Comments on stability of frustrated frustration patterns in vortex language
+In the main text, we pointed out that resistive dips at integer $f+f^{\prime}$ can be stable if $f$ and $f^{\prime}$ are individually close (but not equal) to integer, even though the regular frustration pattern is absent.
+We argued that the increased stability of those features can be understood in terms of loop currents.
+Here we provide an alternative picture based on the more commonly used language of vortices and the Berezinskii–Kosterlitz–Thouless (BKT) transition.
+The energy of the classical 2D Josephson junction array in terms of the number of vortices $v$ per plaquette is given by (see, e.g., Ref.
+[2] for details),
+$H=\sum_{i,j}(v_{i}-f_{i})U_{ij}(v_{j}-f_{j})\ ,$ (S9)
+where the indices $i,j$ address the plaquette positions in 2 dimensions (i.e., similar to the previous section, both $i$ and $j$ implicitly encode a pair of $x$ and $y$ coordinates), and $U_{ij}$ represents the typical logarithmic long-range interaction.
+For regular square lattices $f_{i}=f$ (i.e., all plaquettes experience the same flux), the well-known BKT picture arises.
+At zero flux $f=0$, the zero vortex state $v_{j}=0$ (which is insulating in that no vortices are available to move) corresponds to the ground state, whereas states with finite vortex numbers or vortex pairs (conducting states) are at low temperatures inaccessible due to a finite energy gap, leading to the BKT transition.
+At finite flux, e.g., $f>0$, the system reacts in such a way that the new ground state can be to a good approximation described by a state with a finite vortex density of $\sim f$ vortices per plaquette, in order to screen the applied external flux, where (as is well-known) the propensity of the ground state to have ordered (and thus likewise insulating) vortex configurations or many freely moving vortex configurations depends on whether $f$ is a rational number or not.
+For the checker board flux texture on the other hand, the situation is different in the following crucial aspect.
+The special points of integer $f+f^{\prime}$ can be alternatively regarded as cancelling neighbouring fluxes $f^{\prime}=-f$ (up to modulo one).
+At these special points we still have the zero vortex state $v_{j}=0$ as the gapped ground state because averaged over the whole system, no flux is applied, and hence no finite vortex density occurs due to screening.
+We therefore arrive again at the superconducting BKT phase, notably, at generally irrational values of $f$.
+We emphasise that these zero vortex states are by no means trivial.
+Consider the special (rational) case of $f=1/2$ and $f^{\prime}=-1/2$, with the zero vortex ground state.
+With a simple unitary transformation (increasing the vortex number on every $-1/2$ plaquette by one), we map to the regular square lattice $f=f^{\prime}=1/2$ with the well-known ground state given by an alternating $0,1$ vortex configuration.
+Consequently, $f=-f^{\prime}=1/2$ with the zero vortex ground state and $f=f^{\prime}=1/2$ with alternating $0,1$ vortex ground state are equivalent.
+Now, the difference to the case where $f=-f^{\prime}$ but $f$ irrational (or in fact any rational or irrational number other than $f=0$ or $f=1/2$) is that the correspondence to the regular lattice $f=f^{\prime}$ is no longer exact (as there exists no unitary mapping due to $v_{i}\in\mathbb{Z}$).
+For values close to $1/2$, there still exists a unitary map to a system approximately equivalent to $f=f^{\prime}=1/2$, this correspondence grows weaker and weaker as we move away from $1/2$, which at the value of $1/4$ or lower leads to a cross over, where the checker board system with $f=-f^{\prime}$ more closely imitates the regular $f=f^{\prime}=0$ system instead of $f=f^{\prime}=1/2$.
+Regarding stability, this means that the points of $f+f^{\prime}$ which are closer to $f,f^{\prime}$ integer (than half-integer) have a higher stability, and they can be observed at higher current bias (or higher temperature) before the half integer resistance dips kick in.
+This qualitative picture is in agreement with the findings of the resistively capacitively shunted junction model calculation, see main text, and previous supplementary section.
+SUPPLEMENTAL NOTE 12: Estimation of $\beta$ and the central weak link area
+According to the theoretical discussion in the main text, the parameter $\beta$, the dip periodicity ($p_{\text{dips}}$), and the beating periodicity ($p_{\text{beating}}$), in frustration space, are given by:
+$\beta=\frac{f}{f^{\prime}}=\frac{A}{A^{\prime}},$ (S10)
+$p_{\text{dips}}=\frac{\beta}{\beta+1},$ (S11)
+$p_{\text{beating}}\approx\beta.$ (S12)
+One can use the the difference in periodicities $\Delta p=p_{\text{beating}}-p_{\text{dips}}$ to estimate $\beta$:
+$\Delta p\approx\beta-\frac{\beta}{\beta+1}.$ (S13)
+For our estimate $\Delta p$ was determined by counting the number of resistance dips (10 dips) between the two beating points in Fig. 2 c). This yields $\beta\approx 10.9$.
+The central weak link area of the 4TJJs is connected to the unit cell area of the array ($A_{\text{uc}}$) via $A^{\prime}=A/\beta=(A_{\text{uc}}-A^{\prime})/\beta\Rightarrow$
+$A^{\prime}=\frac{A_{\text{uc}}}{(\beta+1)}.$ (S14)
+Thus, the 4TJJs in the array have an estimated weak link area of $A^{\prime}\approx(165\;\text{nm})^{2}$.
+For comparison, such a central weak link area is sketched on one 4TJJ in Fig.
 S11, assuming the weak link region to be a perfect circle with the corresponding area of $(165\;\text{nm})^{2}$.